### v2

# Dynamic Input Mapping Inversion for Algebraic Loop-Free Control in Hydraulic Actuators

Abstract
The application of nonlinear control schemes to electro-hydraulic actuators often requires several alterations in the design of the controllers during their implementation. This is to overcome the challenges that frequently arise from the inherent complexity of such control algorithms owning to model nonlinearities. Moreover, advanced control solutions for this type of systems often introduce input algebraic loops and chatter, which considerably degrade the tracking performance. This study presents a nonlinear control architecture for hydraulic actuators that comprises low-complexity modules, based on well-established designs that facilitate robust high performance in tracking without introducing the aforementioned limitations. Specifically, the proposed solution consists of two variants of a position controller for the hydraulic cylinder and a dynamic input-mapping inversion module to avoid algebraic loops in the control input. The stability of the closed-loop system is analysed using arguments from Lyapunov theory for cascaded non-autonomous nonlinear systems. The effectiveness of the proposed solution is evaluated on a high-fidelity simulator of a wind turbine pitch system. Appropriate quantitative metrics are finally defined to evaluate the closed-loop system performance in comparison to state-of-the-art nonlinear design.
Index Terms:
electro-hydraulic actuator, nonlinear control, parameter estimation, mapping inversion
I Introduction
Electro-hydrualic actuators (EHA) are commonly used in industrial applications because of their high power density and resilience to harsh environment. Typical applications include aviation, manufacturing, marine/offshore engineering, robotics and energy systems [1, 2, 3].
Although Proportional-Integral-Differential (PID) controllers constitute the industrial standard in these application domains, they often fall short in maintaining high performance in the presence of strong nonlinearities and uncertainties that characterize hydraulic actuator systems. This has motivated the exploration of nonlinear control alternatives that can improve performance and enhance closed-loop robustness over the entire envelope of operation. The complexity, however, of system dynamics often hinders effective application of nonlinear model-based solutions such as feedback linearization (FL) or sliding mode control (SMC) because several limiting assumptions and modelling simplifications need be imposed to apply these methods.
A common heavy simplification concerns the valve dynamics, where two main issues can be identified. A first limitation is encountered when a system model cannot be written as an affine form $\dot{x}=f(x)+g(x)u$, as the vector field of a valve depends non-linearly on the input. A common approach to address this issue is to use the input value at the previous time instant, under the assumption that the input $u$ is slowly varying. This removes the nonlinearity and avoids the introduction of an algebraic loop, at the cost of increased chatter in the control signal [4]. The second design challenge arises with inclusion of actuator dynamics, which increase the relative degree of the system, thus making the design much more complicated, as will be discussed later. Neglecting the valve dynamics during the design stage can simplify implementation but may potentially result in a quite demanding control action going beyond obtainable valve response. This may introduce additional chatter and potential wind up effects.
Furthermore, micro-movements in EHAs are responsible for local heating of hydraulic oil, which degrades oil quality. This leads to increased friction, and eventually to wear and potential damage of a valve [5]. It is therefore essential to establish control designs that combine the performance and robustness features of nonlinear solutions, but avoid excessive actuation demand.
Several studies in the existing literature have reported solutions that adopt nonlinear strategies for controlling hydraulic systems. A cascaded feedback linearization controller with a sliding mode observer was designed in [6]. Sliding mode was explored in a standard fashion in [7], in combination with adaptive control in [8], with an integral sliding surface in [9] and later in [10] to eliminate the velocity state in the control action. Adaptive solutions have been applied by means of backstepping, in e.g. [11], [12], [13], and [14]. The authors in [15] used a coordinate transformation and tuning functions to obtain a backstepping controller that included the valve dynamics. The high relative degree led to a complex solution. Except for the latter, all the other solution adopted the strategy presented in [4] to cancel the algebraic loop, and do not account for the valve dynamics in the design of the controller.
The aim of this paper is to leverage the results on online estimation of nonlinearly parametrised perturbations [16, 17] as a dynamical mapping inversion (MI) algorithm to derive the control input from a virtual input designed via classical nonlinear methods. The paper shows how a modular and straightforward design is achieved, and how the MI can reconstruct the control signal without introducing algebraic loops. Moreover, it is shown how the suggested approach facilitates a systematic commissioning feature, namely, how it can accommodate valve specification parameters such that only frequencies that can actually be delivered by the valve are allowed in the control input. In this way, the performance can be improved and chatter in the input be significantly reduced.
This article contributes by
  1. 1.
    presenting a modular architecture for nonlinear control of hydraulic actuators without introducing algebraic loops or limiting assumptions on the input signal,
  2. 2.
    proposing a scheme for reconstructing the control signal by leveraging a nonlinear parameter estimation algorithm for dynamically inverting the mapping between control input and virtual input,
  3. 3.
    introducing a systematic method to tune the dynamical mapping inversion component according to the actuator parameters,
  4. 4.
    providing rigorous stability and boundedness analyses of the closed-loop dynamics under different control laws,
  5. 5.
    demonstrating the effectiveness of the proposed solution against conventional nonlinear control methods for industrial EHA systems in a real case scenario, using a high-fidelity simulator.
The remainder of the paper is organized as follows. Sec. II presents the hydraulic actuator system and defines the problem at hand. Sec. III designs controllers following commonly accepted standards, and Sec. IV proposes a new algorithm and presents features and benefits in contrast to the standard existing approaches. Stability of the closed-loop system is the subject of Sec. V. Sec. VI presents a case study of a wind turbine pitch control, and results from high fidelity simulation are shown in Sec. VII. Finally, conclusions and future directions are staged in Sec. VIII.
II Problem definition
An electro-hydraulic actuator is composed of an hydraulic cylinder driven by a directional control valve, as shown in Fig. 1. The pressures $p_{p}$ and $p_{r}$ in the piston-side and rod-side chambers of the cylinder are varied according to the valve spool position $x_{v}$. This allows for control of both position and velocity ($x_{c}$ and $v_{c}$) of the the cylinder’s piston. The model for a generic electro-hydraulic actuator is given by [9]:
$M\dot{v}_{c}=A_{p}p_{p}-A_{r}p_{r}-F_{Fr}-F_{ext}$ (1)
$\dot{x_{c}}=v_{c}$
$\dot{p}_{p}=\frac{B_{e,p}(p_{p})}{V_{p}(x_{c})}\left[Q_{p}(x_{v})-A_{p}v_{c}\right]$
$\dot{p}_{r}=\frac{B_{e,r}(p_{r})}{V_{r}(x_{c})}\left[-Q_{r}(x_{v})+A_{r}v_{c}\right]$
$\dot{v}_{v}+2\xi\omega_{0}v_{v}+\omega_{0}^{2}x_{v}=\omega_{0}^{2}u_{x}$
$\dot{x_{v}}=v_{v},$
Figure 1: Generic electro-hydraulic actuator configuration: above, the linear cylinder with piston and rod; below, the flow control valve, oil pump and drain.
where $M$ is the piston mass, $A_{p}$ and $A_{r}$ the cylinder cross sections in the two chambers, $F_{fr}$ and $F_{ext}$ the friction force and the external load respectively, and $B_{e,i}(p_{i})$ the effective bulk modulus of the oil in the $i$-th component. The control input $u_{x}$ is a position reference for the control valve’s spool. The terms
$V_{p}(x_{c})=V_{0,p}+A_{p}x_{c}$ (2)
$V_{r}(x_{v})=V_{0,r}+A_{r}\left(x_{c,m}-x_{c}\right).$
represent the volumes of the cylinder chambers.
The flows are calculated as
$Q_{p}(x_{v})=K(x_{v})\sqrt{\Delta_{pp}(x_{v})}$ (3)
$Q_{r}(x_{v})=K(x_{v})\sqrt{\Delta_{pr}(x_{v})}.$
where the nonlinear function $K(x_{v})$ determine the valve flow coefficient based on its spool position, and
$\Delta_{pp}=(p_{s}-p_{p})\,H\left(x_{v}\right)+(p_{p}-p_{t})\,H\left(-x_{v}\right)$ (4)
$\Delta_{pr}=(p_{r}-p_{t})\,H\left(x_{v}\right)+(p_{s}-p_{r})\,H\left(-x_{v}\right),$
with $H(\cdot)$ being the Heaviside step function.
The control objective in EHA systems is the tracking of a reference signal for the cylinder piston position $x_{c}$. By inspecting the cylinder position dynamics in Eq. 1 it is easily seen that the system corresponding to the position error defined as
$e\triangleq x_{c}-r,$ (5)
has relative degree 5, i.e. $e$ needs to be differentiated five times in order to arrive to a controller canonical form
$e^{(\alpha)}=f(u),$ (6)
which is suitable for control design.
For the class of EHA systems this introduces a high degree of complexity due to the strong nonlinear couplings between the various pressure terms. A common approach adopted in the literature to alleviate this complexity is the introduction of a simplifying assumption, namely the neglection of the valve dynamics. This implies that the position $x_{v}$ inside the terms $Q_{p}(x_{v})$ and $Q_{r}(x_{v})$ can instantaneously assume any desired value and is, therefore, considered the effective control input for the system, i.e. $u\triangleq x_{v}$, with $u\in[-1,1]$.
An additional frequent simplification made in many reported studies is considering one single constant value for the bulk modulus of the oil for both chambers, i.e., $B_{e,p}=B_{e,r}=B>0$. The resulting model, which is most commonly used in the industry for control design, is [18]:
$M\dot{v}_{c}=A_{p}p_{p}-A_{r}p_{r}-F_{Fr}-F_{ext}$ (7)
$\dot{x}_{c}=v_{c}$
$\dot{F}_{h}=A_{p}\dot{p}_{p}-A_{r}\dot{p}_{r}$
$\quad=\frac{A_{p}B}{V_{p}(x_{c})}\big{[}Q_{p}(u)-A_{p}v_{c}\big{]}-\frac{A_{r}B}{V_{r}(x_{c})}\big{[}A_{r}v_{c}-Q_{r}(u)\big{]},$
where the hydraulic force $F_{h}=A_{p}p_{p}-A_{r}p_{r}$ is introduced to avoid considering internal dynamics.
III Baseline nonlinear control design
Considering the model presented in Eq. 7 and the position error $e$ defined in Eq. 5, the dynamics constitutes a relative degree 3 system and reads:
$e^{(3)}=$ $-\frac{\dot{F}_{fric}+\dot{F}_{ext}}{M}-\frac{A_{p}^{2}B}{MV_{p}}v_{c}-\frac{A_{r}^{2}B}{MV_{r}}v_{c}-r^{(3)}+$ (8)
$\frac{A_{p}B}{MV_{p}}Q_{p}(u)+\frac{A_{r}B}{MV_{r}}Q_{r}(u)\triangleq\,\varpi_{FL}.$
Intermediate derivatives are given by
$\dot{e}$ $=v_{c}-\dot{r}$ (9)
$\ddot{e}$ $=-\frac{F_{fric}+F_{ext}}{M}+\frac{A_{p}p_{p}-A_{r}p_{r}}{M}-\ddot{r},$
where the virtual input $\varpi_{FL}=\varpi_{FL}(t,\boldsymbol{x},u)$ is a function of time, of states $\boldsymbol{x}\triangleq\begin{bmatrix}v_{c}&x_{c}&F_{h}\end{bmatrix}^{T}$, and of the control input $u$.
Two basic control solutions are first presented to demonstrate both the effectiveness and the reasons behind the development of a new algorithm. Feedback linearization is presented as a baseline nonlinear control design, which is popular for adaptive nonlinear control solutions. A supertwisting sliding mode control design is shown as a second popular solution.
III-A Feedback linearization (FL)
A FL controller is obtained from writing the error derivative dynamics as follows,
$\dot{\underline{\mathbf{e}}}=\underbrace{\begin{bmatrix}0&1&0\\ 0&0&1\\ 0&0&0\end{bmatrix}}_{\boldsymbol{A_{FL}}}\underbrace{\begin{bmatrix}e\\ \dot{e}\\ \ddot{e}\end{bmatrix}}_{\underline{\mathbf{e}}}+\underbrace{\begin{bmatrix}0\\ 0\\ 1\end{bmatrix}}_{\boldsymbol{B_{FL}}}\varpi_{FL}.$ (10)
Selecting
$\varpi_{FL}=-K_{FL}\underline{\mathbf{e}}$ (11)
with $K_{FL}$ obtained with a standard linear design (eignvalue assignment, LQR, etc.), renders the closed-loop system globally exponentially stable (GES) with the rate of decay being defined by the gain $K_{FL}$. The related control input $u$ is isolated from Eq. 11 taking also Eq. 8, and Eq. 9 into account.
III-B Supertwisting SMC
Sliding mode controllers aim at enforcing the trajectories $\boldsymbol{e}(t)\in\mathbb{R}^{3}$ of the error system onto a pre-designed manifold $\mathcal{S}=\{\boldsymbol{e}\in\mathbb{R}^{3}\,\big{|}\,s(\boldsymbol{e})=0\}$. The variable $s$ is referred to as the sliding variable and is designed to guarantee exponential stability of the error system, so long $s$ is regulated at the origin.
In order to leverage the reduced chatter properties of the well-studied supertwisting sliding mode control designs [19], the dynamics of s should be brought into the form of Eq. 6 with a single differentiation. Therefore, the sliding variable $s$ is defined as:
$s=\ddot{e}+\lambda_{2}\dot{e}+\lambda_{1}e,$ (12)
where $\lambda_{1},\lambda_{2}$ are selected such that the characteristic polynomial in Eq. 12 is Hurwitz. The dynamics of the sliding variable $s$ is:
$\dot{s}=\underbrace{e^{(3)}+\lambda_{2}\ddot{e}+\lambda_{2}\dot{e}}_{\displaystyle\varpi_{ST}(t,\boldsymbol{x},u)},$ (13)
where the right-hand side is of Eq. 13 is again a function $\varpi_{ST}$ of time, the states and the control input $u$. Eq. 12 implies global exponential stability for the origin of the piston position error when the system is in sliding mode, i.e. when $s=\dot{s}=0$. This can be achieved by selecting
$\varpi_{ST}=-k_{1}\lfloor s\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s(\tau))d\tau,$ (14)
where $k_{1},k_{2}>0$. Similar to the controller case in the previous subsection, the control input is derived by isolating the term $u$ in equation 14.
III-C Control input
The final step for designing the controllers is deriving the control input $u$ for which the closed-loop system dynamics is shaped, via the previously presented design methods. From Eq. 8 it is noticed that the error dynamics depends on the terms $F_{ext}$ and $F_{fr}$, and their derivatives. Such terms are in general unknown and need be estimated to compute the control input $u$.
III-C1 Disturbance observer
Although dedicated tests can be run to identify the friction coefficients, it is convenient in this investigation to assume the friction term to be unknown. This allows for lumping it together with the external load in a single term $F_{d}=F_{ext}+F_{fr}$, which can then be estimated with a conventional linear observer. Such an observer can be designed based on the measured pressures as follows,
$\dot{\hat{\boldsymbol{z}}}=\underbrace{\begin{bmatrix}0&-\frac{1}{M}\\ 0&0\end{bmatrix}}_{\boldsymbol{A}}\hat{\boldsymbol{z}}+\begin{bmatrix}\displaystyle\frac{A_{p}p_{p}-A_{r}p_{r}}{M}\\ 0\end{bmatrix}+\underbrace{\begin{bmatrix}L_{1}\\ L_{2}\end{bmatrix}}_{\boldsymbol{L}}\underbrace{\begin{bmatrix}1&0\end{bmatrix}}_{\boldsymbol{C}}\tilde{\boldsymbol{z}},$ (15)
where $\hat{\boldsymbol{z}}=\begin{bmatrix}\hat{v_{c}}&\hat{F}_{d}\end{bmatrix}^{T}$, $\tilde{\boldsymbol{z}}\triangleq\boldsymbol{z}-\hat{\boldsymbol{z}}$ and $\boldsymbol{L}$ chosen such that $\boldsymbol{A}-\boldsymbol{LC}$ is Hurwitz, ensuring that the disturbance term $F_{d}$ can be reconstructed with a bounded error [20]. Furthermore, since $\boldsymbol{z}\in\mathcal{L}_{\infty}$ and $\boldsymbol{A}-\boldsymbol{LC}$ is Hurwitz by design, it holds that
$\|\boldsymbol{\dot{\tilde{z}}}\|=\|(\boldsymbol{A}-\boldsymbol{LC})\,\boldsymbol{\tilde{z}}\|\leq-\lambda_{min}(\boldsymbol{A}-\boldsymbol{LC})\,\|\boldsymbol{\tilde{z}}\|_{\infty}\triangleq L_{z},$ (16)
where $L_{z}$ is a positive real number. Thus, $\dot{F}_{d}$ is also reconstructed with a bounded error. Global asymptotic stability of the entire closed-loop system will be proved subsequently.
III-C2 Control input derivation
Directly isolating $u$ in Eq. 11 and Eq. 13 is not possible: while it may be possible to invert the $K(u)$ function in Eq. 3, the control input also appears inside the Heaviside functions, generating an algebraic loop where the derived input depends on the input itself. The standard way to solve the problem is the one described in [4]:
$Q_{p}(t,u)=$ $K(u)\left[\sqrt{p_{s}-p_{p}}\,H\left(u\right)+\sqrt{p_{p}-p_{t}}\,H\left(-u\right)\right]$ (17)
$\approx$ $Ku\left[\sqrt{p_{s}-p_{p}}\,H\left(u_{-1}\right)+\sqrt{p_{p}-p_{t}}\,H\left(-u_{-1}\right)\right]$
$=$ $K_{Qp}\,u,$
$Q_{r}(t,u)=$ $K(u)\left[\sqrt{p_{r}-p_{t}}\,H\left(u\right)+\sqrt{p_{s}-p_{r}}\,H\left(-u\right)\right]$ (18)
$\approx$ $Ku\left[\sqrt{p_{r}-p_{t}}\,H\left(u_{-1}\right)+\sqrt{p_{s}-p_{r}}\,H\left(-u_{-1}\right)\right]$
$=$ $K_{Qr}\,u,$
where the term $u_{-1}$ represents the value of the input at the previous time step. In this way, the algebraic loop is eliminated and $u$ enters linearly in both equations. By defining
$g=\frac{A_{p}B}{MV_{p}}K_{Qp}+\frac{A_{r}B}{MV_{r}}K_{Qr},$ (19)
the standard feedback linearization controller is then,
$u_{FL}=\frac{1}{g}\left[\frac{\dot{\hat{F}}_{d}}{M}+\frac{A_{p}^{2}B}{MV_{p}}v_{c}+\frac{A_{r}^{2}B}{MV_{r}}v_{c}+r^{(3)}-K_{FL}\underline{\mathbf{e}}\right],$ (20)
and the standard super-twisting sliding mode controller is,
$u_{ST}=\frac{1}{g}$ $\left[\frac{\dot{\hat{F}}_{d}}{M}+\frac{A_{p}^{2}B}{MV_{p}}v_{c}+\frac{A_{r}^{2}B}{MV_{r}}v_{c}-\lambda_{2}\ddot{e}-\lambda_{1}\dot{e}\right.$ (21)
$\,\,\left.+r^{(3)}-k_{1}\lfloor s\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s)d\tau\right].$
The $K(u)$ linearization is then reflected into uncertainties in the input term $g$ and in the terms $K_{Qp}$ and $K_{Qr}$.
IV Cascaded control design
The delay operator introduced to avoid the formation of algebraic loops simplified the controller implementation with the drawback of increasing unwanted chatter on the control signal, with the consequences on oil quality described above. The solution proposed in this article for avoiding this, is to leverage a virtual input for the controller design and dynamically deriving a real input that produces the virtual one required by the controller. The block diagram is shown in Fig. 2.
Figure 2: Block diagram of the proposed algorithm: Cascaded control with dynamic input mapping inversion.
By defining a virtual input as
$v\triangleq\frac{A_{p}B}{MV_{p}}Q_{p}(u)+\frac{A_{r}B}{MV_{r}}Q_{r}(u),$ (22)
Eq. 8 can be rewritten as
$e^{(3)}=-\frac{\dot{F}_{fric}+\dot{F}_{ext}}{M}-\frac{A_{p}^{2}B}{MV_{p}}v_{c}-\frac{A_{r}^{2}B}{MV_{r}}v_{c}-r^{(3)}+v,$ (23)
where the virtual input $v$ is a function of time and the control input $u$, and is now present as a linear term. The feedback linearizing and supertwisting control designs are now defined in terms of the associated virtual inputs as, respectively,
$v_{FL}=\frac{\dot{\hat{F}}_{d}}{M}+\frac{A_{p}^{2}B}{MV_{p}}v_{c}+\frac{A_{r}^{2}B}{MV_{r}}v_{c}+r^{(3)}-K_{FL}\underline{\mathbf{e}},$ (24)
and
$v_{ST}=$ $\frac{\dot{\hat{F}}_{d}}{M}+\frac{A_{p}^{2}B}{MV_{p}}v_{c}+\frac{A_{r}^{2}B}{MV_{r}}v_{c}-\lambda_{2}\ddot{e}-\lambda_{1}\dot{e}$ (25)
$+r^{(3)}-k_{1}\lfloor s\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s(\tau))\,d\tau.$
IV-A Dynamic mapping inversion algorithm
The last part of the proposed method constitutes the derivation of the control input $u$ by solving Eq. 22 for $u$ in each case. A dynamic nonlinear mapping inversion scheme is adopted from [16] for this purpose, allowing for re-formulating of the input calculation into a parameter estimation problem in $u$ for a nonlinearly parametrized regressor.
Specifically, let $\rho:[0,\infty)\times\mathbb{R}^{3}\times[-1,1]\rightarrow\mathbb{R}$ with
$\rho(t,\boldsymbol{x},u)=\frac{A_{p}B}{MV_{p}}Q_{p}(t,\boldsymbol{x},u)+\frac{A_{r}B}{MV_{r}}Q_{r}(t,\boldsymbol{x},u)$ (26)
be a mapping from any given actual control input $u$ to the virtual input that was earlier defined in Eq. 22. Assume that for every value $v$ at each time instant there exists an appropriate value of the control input $u^{*}$ such that $v=\rho(u^{*})$. It has been shown [16, 17] that the estimation law
$\dot{u}=\gamma\,\mu(t,\boldsymbol{x},u)\,\Big{(}v-\rho(t,\boldsymbol{x},u)\Big{)},$ (27)
where $\gamma>0$ is a tuning gain and $\mu(t,\boldsymbol{x},u)$ a function to be designed, ensures exponential convergence of the estimation error $\tilde{u}\triangleq u^{*}-u$ to the origin if the following conditions hold for all pairs $(u^{*},u)$:
Condition 1.
There exists a positive definite real function $\sigma(t,\boldsymbol{x})$ such that
$\mu(t,\boldsymbol{x},u^{*})\left.\frac{\partial\rho}{\partial u}\right|_{u}+\left.\frac{\partial\rho}{\partial u}\right|_{u^{*}}\mu(t,\boldsymbol{x},u)\geq 2\sigma(t,\boldsymbol{x})~{}.$ (28)
Condition 2.
There exists a positive real number $L_{\mu}>0$ such that
$\left|\rho(t,\boldsymbol{x},u^{*})-\rho(t,\boldsymbol{x},u)\right|\leq L_{\mu}|\tilde{u}|\sqrt{\sigma(t,\boldsymbol{x})}\;.$ (29)
Condition 3 (Persistency of excitation).
There exist positive constants $T,\alpha_{0}>0$ such that
[…]
where $\boldsymbol{I}$ is the identity matrix.
Remark 1.
If the mapping $\rho$ is not injective, then under Conditions 1-3, the estimation law in Eq. 27 will ensure that $u$ converges to an appropriate value such that $\rho(t,u)$ converges to $v$ exponentially fast.
Figure 3: Lookup table approximation with polynomial functions.
A typical choice for $\mu$ is the gradient of $\rho$ with respect to $u$. However, computation of $\mu$ requires differentiation of the nonlinear function $K(u)$ present in Eq. 3 and, since the latter is usually available as a lookup table, it is approximated by an appropriate polynomial function of $u$. Fig. 4 illustrates an example of different polynomial approximations for $K(u)$, i.e.
$K(u)=\sum_{i=0}^{n}\alpha_{i}\,u^{i},$ (31)
with a fifth-order polynomial yielding the best fit.
This particular approximation also ensures that $\frac{\partial K}{\partial u}>~{}0,\;\forall u\in[-1,1]$ as shown in Fig. 4, which will be useful later in proving the validity of the convergence conditions. The function of $\mu$ can now be selected as
$\mu(t,\boldsymbol{x},u)\triangleq\frac{\partial K}{\partial u}\left(\frac{A_{p}B}{MV_{p}}\sqrt{\Delta_{pp}(u)}+\frac{A_{r}B}{MV_{r}}\sqrt{\Delta_{pr}(u)}\right)>0,$ (32)
since the quantity in the parenthesis is always positive. Evaluation of Condition 1 leads to
$\mu(t,\boldsymbol{x},u^{*})\left.\frac{\partial\rho}{\partial u}\right|_{u}$ $+\mu(t,\boldsymbol{x},u)\left.\frac{\partial\rho}{\partial u}\right|_{u^{*}}=2\mu(t,\boldsymbol{x},u^{*})\mu(t,\boldsymbol{x},u)$ (33)
$\geq 2\left(\inf\limits_{u}\mu(t,\boldsymbol{x},u)\right)^{2}\triangleq\sigma(t,\boldsymbol{x})>0\;.$
Moreover, since $\frac{\partial\rho}{\partial u}$ is bounded, there exists a Lipschitz constant $L_{u}>0$ such that
$\left|\rho(t,\boldsymbol{x},u^{*})-\rho(t,\boldsymbol{x},u)\right|<L_{u}|\tilde{u}|\leq\frac{L_{u}|\tilde{u}|\sqrt{\sigma(t,\boldsymbol{x})}}{\sqrt{\inf\limits_{(t,\boldsymbol{x})}\sigma(t,\boldsymbol{x})}}$ (34)
which fulfills Conditions 1,2 with
$L_{\mu}=\displaystyle\frac{L_{u}}{\sqrt{\inf\limits_{(t,\boldsymbol{x})}\sigma(t,\boldsymbol{x})}}\;.$
Finally, Condition 3 is satisfied by the definition of $\sigma$ since $\sigma(t,\boldsymbol{x})>0$, $\forall(t,\boldsymbol{x})\in\mathbb{R}_{\geq 0}\times\mathbb{R}^{3}$.
Remark 2.
The $\mu$ function that was chosen for the purpose is an approximation of the gradient, where the dependency on u in the terms $\Delta_{pp}$ and $\Delta_{pr}$ is neglected during differentiation, since it only appears inside a Heaviside function. Hence, it would only be reflected in additional spikes in case a discontinuity is present during switching.
Figure 4: Verification of derivative for algorithm convergence condition.
V Stability Analysis
The stability of the closed-loop system is analyzed by including the nonlinear controller, the dynamical mapping inversion, and the disturbance estimator. By writing it in the form of a feedback interconnected system
[…]
and by fulfilling the following conditions:
Condition 4.
Unperturbed $\Sigma_{1}$ $(\tilde{u}=0)$ is Uniformly GES
Condition 5.
Solutions of $\Sigma_{2}$ are Uniformly Ultimately Bounded (UUB)
Condition 6.
$\forall(t,\boldsymbol{\zeta})\in[0,\infty)\times\mathbf{R}^{2}$
$\frac{\partial V(t,\boldsymbol{\zeta})}{\partial\boldsymbol{\zeta}}h(t,\tilde{u})=o(W(\boldsymbol{\zeta})),$ (36)
the overall closed-loop system is globally UUB according to Theorem 1 from [21].
Let $v^{*}=\rho(u^{*})$ be the required virtual controller and $u^{*}$ the related control input, one can define
$\tilde{v}\triangleq v^{*}-v=\rho(u^{*})-\rho(u),$ (37)
for some arbitrary $u$ and $v$, with $\tilde{u}=u^{*}-u$ having been defined earlier. The two closed-loop systems obtained by closing the loop with the FL controller and the ST-SMC controller are, respectively:
$\Sigma_{1,FL}:\dot{\underline{\mathbf{e}}}=(\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}}K_{fl})\,\underline{\mathbf{e}}+\boldsymbol{\xi_{F_{d}}}+\boldsymbol{h_{FL}}(t,\tilde{u})$ (38)
$\Sigma_{2,FL}:\dot{\tilde{u}}=\dot{u}^{*}-\gamma\,\mu(x,u^{*},\tilde{u})\,\Big{(}-K_{FL}\underline{\mathbf{e}}-\rho(u^{*}-\tilde{u})\Big{)}$
with
$\boldsymbol{\xi}_{F_{d}}=\boldsymbol{B_{FL}}\left(-\frac{\dot{\tilde{F}}_{d}}{M}+K_{FL}\begin{bmatrix}0\\ -\displaystyle\frac{\tilde{F}_{d}}{M}\end{bmatrix}\right)$ (39)
$\boldsymbol{h_{FL}}(t,\tilde{u})=-\boldsymbol{B_{FL}}\,\rho(\tilde{u}),$
and
$\Sigma_{1,ST}:\dot{\boldsymbol{\vartheta}}=f_{\boldsymbol{\vartheta}}(t,\boldsymbol{\vartheta})+\boldsymbol{h_{ST}}(t,\tilde{u})$ (40)
$\Sigma_{2,ST}:\dot{\tilde{u}}=\dot{u}^{*}-\gamma\,\mu(x,u^{*},\tilde{u})\,\Big{(}v^{*}(t,\boldsymbol{\vartheta})-\rho(u^{*}-\tilde{u})\Big{)}$
with
$f_{\boldsymbol{\vartheta}}(t,\boldsymbol{\vartheta})$ $=\frac{\partial T}{\partial\boldsymbol{\zeta}}\left\{\frac{1}{|\zeta_{1}|}\begin{bmatrix}-\frac{k_{1}}{2}&1\\ -k_{2}&0\end{bmatrix}T^{-1}(\boldsymbol{\vartheta})-\begin{bmatrix}0\\ d\end{bmatrix}\right\}$ (41)
$\boldsymbol{h_{ST}}(t,\tilde{u})$ $=-\frac{\partial T}{\partial\boldsymbol{\zeta}}\begin{bmatrix}\displaystyle\frac{1}{|\zeta_{1}|}\\ 0\end{bmatrix}\rho(\tilde{u}-\hat{u})$
according to [22], where $v^{*}(t,\boldsymbol{\zeta})$ was defined in Eq. 25 and $\boldsymbol{\zeta}$ corresponds to a coordinates transformation of the closed-loop dynamics. The detailed derivations are presented in Appendix.
Condition 4 coincides with the proof of stability of the controllers in their standard configuration, and is ensured by choosing the Lyapunov functions as:
$V_{FL}$ $=\underline{\mathbf{e}}^{T}P_{FL}\,\underline{\mathbf{e}}$ (42)
$V_{ST}$ $=\boldsymbol{\vartheta}^{T}P_{ST}\,\boldsymbol{\vartheta}.$
Condition 5 is satisfied according to the result in [23], where it was proven that if $|\dot{u}^{*}|<\infty\,\forall t>0$, then $\tilde{u}(t)$ is UUB for all time.
Finally, Condition 6 is verified as, $\Sigma_{1,FL}$:
$\left\|\frac{\partial V_{FL}}{\partial\underline{\mathbf{e}}}\boldsymbol{h_{FL}}(t,\tilde{u})\right\|$ $\leq\left\|2P_{FL}\underline{\mathbf{e}}\right\|\,\left\|B_{FL}\right\|\,\left|\rho(u^{*})-\rho(u^{*}-\tilde{u})\right|$ (43)
$\leq\left\|2P_{FL}\underline{\mathbf{e}}\right\|\,\left\|B_{FL}\right\|\,L_{p}\,\left\|\tilde{u}\right\|$
$\leq\underbrace{\left\|2P_{FL}\right\|\,\left\|B_{FL}\right\|\,L_{p}\,\left\|\tilde{u}\right\|_{\infty}}_{\eta}\,\left\|\underline{\mathbf{e}}\right\|,$
and for $\Sigma_{1,ST}$,
$\left\|\frac{\partial V_{ST}}{\partial\boldsymbol{\vartheta}}\boldsymbol{h_{ST}}(t,\tilde{u})\right\|$ $\leq\left\|2P_{ST}\boldsymbol{\vartheta}\right\|\frac{L_{\rho}}{\min_{t}(|\zeta_{1}(t)|)}\|\tilde{u}\|_{\infty}\left\|\frac{\partial\tau}{\partial\boldsymbol{\xi}}\right\|_{\infty}$ (44)
$\leq\underbrace{\left\|2P_{ST}\right\|\frac{L_{\rho}}{\min_{t}(|\zeta_{1}(t)|)}\|\tilde{u}\|_{\infty}\left\|\frac{\partial\tau}{\partial\boldsymbol{\xi}}\right\|_{\infty}}_{\eta^{\prime}}\left\|\boldsymbol{\vartheta}\right\|.$
VI Case Study
The case study on which the algorithm is presented is a wind turbine pitch system with the hydraulic actuator used to provide a time-varying blade pitch, mainly according to prevailing wind and power demand. The mechanical link between cylinder and turbine blade is stiff, such that controlling the blade angle is equivalent to controlling cylinder position. The angular dynamics of the blade is described by:
$J_{tot}\dot{v_{\beta}}=\left(A_{p}p_{p}-A_{r}p_{r}\right)r\sin\left(\psi\left(x_{c}\right)\right)-M_{fric}-M_{ext}$ (45)
$\dot{\beta}=v_{\beta},$
where the equivalent mass $M$ also accounts for blade inertia. It is noticed that the equations match the dynamics of the cylinder dynamics presented in Eq. 7. The the presence of a cyclic component in the variable distance that relates mechanical moment to force is minor, as will be shown. Thus, the control design aims at tracking the angle while the cylinder dynamics represents a stable internal dynamics, under the assumption of having a stiff mechanical connection between cylinder and blade. Hence, there exists a static transformation relating $x_{c}$ to $\beta$ and with the pitch control error being the difference between pitch angle and a reference,
$e=\beta-r,$ (46)
the relative degree of the closed loop pitch control will be the same as for the closed loop cylinder piston position control loop analysed above. With relative degree three of both, the third derivative of the error is needed,
$e^{(3)}=$ $-\frac{\dot{M}_{fric}+\dot{M}_{ext}}{J_{tot}}$ (47)
$-\frac{lA_{p}^{2}B}{J_{tot}V_{p}}v_{c}\sin{\psi_{p}(x_{c})}-\frac{lA_{r}^{2}B}{J_{tot}V_{r}}v_{c}\sin{\psi_{p}(x_{c})}$
$+\frac{lA_{p}B}{J_{tot}V_{p}}Q_{p}(u)\sin{\psi_{p}(x_{c})}+\frac{lA_{r}B}{J_{tot}V_{r}}Q_{r}(u)\sin{\psi_{p}(x_{c})}$
$+\underbrace{\frac{l}{J_{tot}}F_{h}\cos(\psi_{p}(x_{c}))\frac{d\psi}{dx_{c}}v_{c}}_{\tau(x)}-r^{(3)}\,\triangleq\,\varpi.$ (48)
With respect to the general case presented in Eq. 8, two variations are present:
  • $1/M$ is replaced by $l\sin{\psi_{p}(x_{c})}/J_{tot}$.
  • An additional term $\tau(x)$, is introduced.
These changes arise due to the variable-length moment arm in the updated dynamics.
Figure 5: Momentum arm angle for verifying that constant assumption holds.
However, Fig. 5 suggests that $\psi(x_{c})$ is approximately constant, and thus the following assumptions are made:
  • $\psi(x_{c})\approx\textit{const}$
  • $\frac{d\psi}{dx_{c}}\approx 0\rightarrow\tau(x)\approx 0$.
VI-A Standard controllers for study case
The control action for the standard cases are obtained by following the procedure described in Sec. III. The input gain is defined as
$g=\frac{lA_{p}B}{J_{tot}V_{p}}K_{Qp}+\frac{lA_{r}B}{J_{tot}V_{r}}K_{Qr}.$ (49)
The feedback linearization controller is derived as,
[…]
where the angle dependency is removed, since it only affects the $r^{(3)}$ term and the term associated with the gain, which is a degree of freedom anyway.
The same applies for the sliding mode controller:
$u_{ST}=\frac{1}{g}$ $\left[\frac{\dot{\hat{M}}_{d}}{J_{tot}}+\frac{lA_{p}^{2}B}{J_{tot}V_{p}}v_{c}+\frac{lA_{r}^{2}B}{J_{tot}V_{r}}v_{c}+r^{(3)}-\lambda_{2}\ddot{e}-\lambda_{1}\dot{e}\right.$ (51)
$\,\,\left.-k_{1}\lfloor s\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s)d\tau\right],$
with
$\begin{bmatrix}\dot{e}\\ \ddot{e}\end{bmatrix}\approx\begin{bmatrix}v_{\beta}-\dot{r}\\ \displaystyle-\frac{\hat{M_{d}}}{J_{tot}}+\frac{l(A_{p}p_{p}-A_{r}p_{r})}{J_{tot}}-\ddot{r}\end{bmatrix}$ (52)
and $\hat{M_{d}}$ the new disturbance being estimated.
VI-B Mapping inversion based controllers for study case
Following the same approach, the virtual control input in the proposed solutions are:
$v_{FL}=\frac{\dot{\hat{M}}_{d}}{J_{tot}}+\frac{lA_{p}^{2}B}{J_{tot}V_{p}}v_{c}+\frac{lA_{r}^{2}B}{J_{tot}V_{r}}v_{c}+r^{(3)}-K_{FL}\underline{\mathbf{e}}$ (53)
$v_{ST}=$ $\frac{\dot{\hat{M}}_{d}}{J_{tot}}+\frac{lA_{p}^{2}B}{J_{tot}V_{p}}v_{c}+\frac{lA_{r}^{2}B}{J_{tot}V_{r}}v_{c}+r^{(3)}-\lambda_{2}\ddot{e}-\lambda_{1}\dot{e}$ (54)
$\,\,-k_{1}\lfloor s\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s(\tau))d\tau,$
and the dynamical mapping inversion algorithm is,
$\dot{u}=\gamma\,\mu(t,\boldsymbol{x},u)\big{(}v-\rho(t,\boldsymbol{x},u)\big{)}$ (55)
$\rho(t,\boldsymbol{x},u)=\frac{lA_{p}B}{J_{tot}V_{p}}Q_{p}(p_{p},p_{r},u)+\frac{lA_{r}B}{J_{tot}V_{r}}Q_{rv}(p_{r},u)$
$\mu(t,\boldsymbol{x},u)=\frac{\partial K}{\partial u}\left(\frac{A_{p}B}{MV_{p}}\sqrt{\Delta_{pp}(u)}+\frac{A_{r}B}{MV_{r}}\sqrt{\Delta_{pr}(u)}\right),$
where $K(u)$ is defined according to Eq. 31 as
$K(u)=10^{-7}\left(-23.67u^{5}+40.73u^{3}+1.97u\right)\;.$ (56)
VII Simulation Results
The designed control actions are tested on a high-fidelity simulator, which includes the general dynamics presented in Eqs. 1,45, and also encompass the control valve dynamics. The measured signals are subjected to noise, simulated at realistic levels. The velocities ($v_{c}$ and $v_{\beta}$) are obtained via numerical differentiation since they are not usually available as sensor signals in a real plant. The pitch angle reference and the disturbance $M_{ext}$ are obtained from a high-fidelity simulator for a wind turbine based on OpenFAST that is widely used in wind turbine research and development.
Figure 6: Tracking performance of the Feedback linearization based controller, with and without applying the dynamical mapping inversion.
VII-A Tracking performance
The tracking performance of the feedback linearization controller is presented in Fig. 6, and in Fig. 7 in a zoomed version. Three controllers are presented, for comparing the standard Feedback Linearizatrion (FL) solution with the one where the dynamic mapping inversion is applied (MI-FL) under the same tuning gains, and by overtuning (OT) the standard solution to be able to match the performance of the MI-FL controller. While both controllers are showing good tracking performance, the MI-FL (for a given tuning) shows a smaller error over the simulation window. By increasing the gains, it is possible to match its performance, with the consequence of getting an increase of unwanted chattering in the input.
Figure 7: Zoomed on regions of interest for showing the tracking properties under different conditions.
The zoomed plot gives a visualization of the input behaviour $u$, and of the valve position $x_{v}$. The MI-FL controller shows a reduced chatter in the regions where the pitch is varying in amplituide, while it shows a more chattery behaviour when the pitch is constant or slowly varying. Furthermore, the signal quality in $u$ and $x_{v}$ are equivalent for the MI solution, while part of the noise is filtered out by the valve dynamics in the standard solution, since the controller doesn’t have knowledge of the valve dynamics, nor of valve limitations.
The results for the ST-SMC are stated in Fig. 8.
Figure 8: Tracking performance of the Supertwisting sliding mode controllers, with and without applying dynamical mapping inversion.
The tracking performances are more similar between the two solutions with respect to the FL controller. The OT version doesn’t improve the tracking performance although leading to a more chattery input. The disturbance rate that the controller needs to handle, is shown in Fig. 9, where it is evident how higher gains lead to a larger amplitude in the command signal.
Figure 9: Perturbation derivative of sliding mode dynamics under different STSMC.
The unexpected dependency of the disturbance rate on the tuning gains is justified by the use of an observer for estimating the disturbance. Higher gains lead to a higher error in the estimation, which is affecting the overall control performance. The performance comparison of the observer with the different controllers is stated in Fig. 10. The tuning doesn’t coincide with the bounds since higher gains are reflected with higher disturbance estimation errors and, thus, higher disturbance derivative bounds. The STSMC is not capable of ensuring convergence on the origin at any time instant, but boundedness is guaranteed, according to [24].
VII-B Mapping inversion algorithm
The tracking performance of the mapping inversion algorithm is presented in Fig. 11. $v$ and $\rho$ are the signal defined in Eq. 55, while $v_{sim}$ is the equivalent of the $v$ signal which is obtained in simulation, hence based on the real states of the system and not on the noisy measurements. The algorithm is not able to perfectly reconstruct $v$, because of the presence of noise in the signals. However, it is shown how it is possible to determine a proper tuning such that it is possible to reconstruct the actual virtual input in the plant ($v_{sim}$) and filtering out the noise present in the measurements.
Figure 10: Wind load disturbance estimation and comparison with LPF hydraulic force.
Thus, there is a natural filtering action which can be approximate with a simple first order low-pass filter in order to determine an indication of the bandwidth of such filtering action, in comparison to the bandwidth of the valve which represents a physical limitation in what the control action can deliver on the real system. The comparison is presented in Fig. 12.
Figure 11: Dynamical mapping inversion algorithm capability of tracking the required virtual controller, with a filtering action that allows to reconstruct the actual signal value in simulation
The cutoff frequency of the MI algorithm is lower than the valve’s bandwidth, meaning the valve only needs to follow trajectories it can actually withstand. With respect to the standard solution, where the filtering was occurring in the valve directly, in this new solution is the algorithm accounting for the valve limitations and filtering the input signal. This result support the similar signal quality in $u$ and $x_{v}$ in the MI solution that was found in Fig. 7.
Figure 12: Bode plots of valve dynamics and mapping inversion filtering effect.
Remark 3.
The tuning presented shows some room with respect to the maximum bandwidth that can be achieved by the valve. However, a faster tuning is reflected in a noisier reconstruction of $v_{sim}$ from the noisy measurements, hence more chattery $u$ and $x_{v}$. The slew rate limitation of the valve may also play a role.
VII-C Controllers comparison
While the plots are able to display some clear properties of the system and some qualitative performance indications can be obtained from them, a more rigorous analysis is performed to have a quantitative description and comparison of the proposed solutions. The controllers are tested according to some metrics, related both to the tracking performance and to the input cleanliness.
Assessing the tracking performance based on the error is straightforward, while assessing the signal quality of the input is more challenging since it is not a constant average signal. The solution was to calculate a moving average $\bar{u}$ and defining a new signal $u_{f}=u-\bar{u}$ which has zero mean and still maintains the information of the high frequency oscillation. The moving average window is the same for all the tests. The same applies for determining $x_{v,f}$. An example of a $u_{f}$ signal is illustrated in Fig. 13.
The following metrics are defined for the purpose:
  • $\text{RMS}(u_{f})=\sqrt{\frac{1}{N}\sum_{i=1}^{N}u_{fi}^{2}}$
  • $\text{RMS}(x_{v,f})=\sqrt{\frac{1}{N}\sum_{i=1}^{N}x_{v,fi}^{2}}$
  • $\text{MAX}(e)=\max\limits_{0\leq t\leq T}e(t)$
  • $\text{RMS}(e)=\sqrt{\frac{1}{N}\sum_{i=1}^{N}e_{i}^{2}}$
  • $\text{QWAE}(e)=\sqrt{\frac{\sum_{i=1}^{N}e_{i}^{2}|e_{i}|}{\sum_{i=1}^{N}e_{i}^{2}}}$
[…]
where $T$ is the simulation time and the resulting values are normalized for having a clearer comparison.
Figure 13: Moving average removal for deriving a metrics associated with signal chatter.
The QWAE (Quadratically Weighted Average Error) metric is used to penalize larger errors more heavily, while the ISE (Integral Squared Error) metric measures the persistence of errors.
The result for the system running with the different controllers is shown in Fig. 14.
The solution proposed in the article performs better than standard one with the same tuning and with higher gains. The only exceptions are the cases of the $RMS(e)$ and the $ISE(e)$, where the OT ST-SMC displays a slightly smaller error, at the price of having much worse input performance. For a fixed tuning, the error with the MI is lower than when applying the respective standard algorithm.
The input signal is significantly improved by the filtering action of the MI, while the improvement in the actual valve position becomes less significant thanks to the filtering action of the valve itself. Still, the solution that leverages the MI shows the cleanest input over a window where both fast varying and constant references are required. The amplitude of $\text{RMS}(u_{f})$ and $\text{RMS}(x_{v,f})$ are equivalent in the MI case.
Figure 14: Bar plots for evaluating performance of the different controllers according to different metrics, associated with control input and tracking error.
VIII Conclusion
This article presented a solution to eliminate the formation of an algebraic loop when designing some conventional nonlinear controllers. This was achieved by design of a virtual input and dynamically deriving a control action, which emulated the desired behaviour, while control valve trajectory commands were avoided if they would exceed the specified admissible range for valve behaviours.
The effectiveness of the proposed solution was tested in high-fidelity simulation for the industrial application of wind turbines’ blade pitch control, showing a general improvement in tracking performance and an essential reduction of input chatter, especially in the regions where the pitch reference is highly varying.
The dynamic inversion solution enabled room for flexibility and allows for future progress with other nonlinear parameter estimation algorithms.
A further step towards industrial implementation will be to derive the discrete-time version of the algorithm.
Acknowledgements
This research was funded by The Danish Energy Technology Development and Demonstration Programm (EUDP) through the project: “Decreased Cost of Energy (CoE) from wind turbines by reducing pitch system faults”, grant number 64022-1058. The authors much appreciate this support.
[Feedback interconnected systems derivations] Let $\delta_{\underline{\mathbf{e}}}$ be the error associated with the calculation of $\ddot{e}$. The closed-loop system with the FL controller in the form of $\Sigma_{FL}$ is obtained as:
$\dot{\underline{\mathbf{e}}}=$ $\boldsymbol{A_{FL}}\underline{\mathbf{e}}+\boldsymbol{B_{FL}}(v^{*}-\tilde{v})$ (57)
$=$ $\boldsymbol{A_{FL}}\underline{\mathbf{e}}+\boldsymbol{B_{FL}}\left(-K_{FL}(\underline{\mathbf{e}}-\delta_{\underline{\mathbf{e}}})-\frac{\dot{\tilde{F}}_{d}}{M}-\tilde{v}\right)$
$=$ $(\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}}K_{fl})\,\underline{\mathbf{e}}+\boldsymbol{B_{FL}}\left(K_{FL}\begin{bmatrix}0\\ -\frac{\tilde{F}_{d}}{M}\end{bmatrix}-\frac{\dot{\tilde{F}}_{d}}{M}\right)$
$-\boldsymbol{B_{FL}}\,\left[\rho(\tilde{u})-\rho(u^{*}-\tilde{u})\right]$
$=$ $(\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}}K_{fl})\,\underline{\mathbf{e}}+\boldsymbol{\xi_{F_{d}}}+\boldsymbol{h_{FL}}(t,\tilde{u}).$
The unperturbed system is $\boldsymbol{h_{FL}}(t,0)=\boldsymbol{0}$ and the remaining terms describe the closed-loop dynamics with the standard design. Except for the bounded term $\boldsymbol{\xi_{F_{d}}}$, the system is exponentially stable with a convergence rate determined by the eigenvalues of $\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}}K_{fl}$. The bounded term acts as a disturbance that causes the system to be UUGB with sufficiently high gains.
Let
$\tilde{s}=s-s^{*},$ (58)
where $s$ is the actual value of the sliding variable and $s^{*}$ is calculated by differentiating the error with a robust exact differentiator, like the one in [25]. The sliding variable dynamics in closed loop reads:
$\dot{s}=$ $-k_{1}\lfloor s^{*}\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s^{*}(\tau))d\tau$ (59)
$-k_{1}\lfloor\tilde{s}\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(\tilde{s}(\tau))d\tau-\frac{\dot{\tilde{F}}_{d}}{M}-\lambda_{2}\frac{\tilde{F}_{d}}{M}$
$-\rho(u^{*}-\tilde{u})$
$=$ $-k_{1}\lfloor s^{*}\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s^{*}(\tau))d\tau-\Delta(t)-\rho(u^{*}-\tilde{u}),$
where all the terms in $\Delta$ are bounded: $\tilde{s}$ is bounded and with zero mean, since the reconstruction error with the robust exact differentiator is only related to noise, and the disturbance estimations are bounded (proof in Sec. III-C1). For $s\neq 0$ one can define [22]
$\zeta_{1}\triangleq|s^{*}|^{1/2}sgn(s^{*})$ (60)
[…]
$\boldsymbol{\zeta}\triangleq\begin{bmatrix}\zeta_{1}&\zeta_{2}\end{bmatrix}^{T},$
and rewrite the closed-loop system as:
$\boldsymbol{\dot{\zeta}}=\frac{1}{|\zeta_{1}|}\begin{bmatrix}-\displaystyle\frac{1}{2}k_{1}&1\\ -k_{2}&0\end{bmatrix}\boldsymbol{\zeta}+\begin{bmatrix}0\\ -d\end{bmatrix}+\underbrace{\begin{bmatrix}-\displaystyle\frac{\rho(u-\tilde{u})}{|\zeta_{1}|}\\ 0\end{bmatrix}}_{\displaystyle\boldsymbol{h}(t,\tilde{u})},$ (61)
with $k_{1},k_{2}$ selected to fulfil the conditions in [26]. Since $h(t,0)=0$, the unperturbed system has a finite stable equilibrium point at the origin. This implies (see [22]) the existence of a diffeomorphism $T:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ such that for $\boldsymbol{\vartheta}=T(\boldsymbol{\zeta})$ and
$\boldsymbol{\dot{\vartheta}}=\frac{\partial T}{\partial\boldsymbol{\zeta}}\left\{\frac{1}{|\zeta_{1}|}\begin{bmatrix}-\frac{k_{1}}{2}&1\\ -k_{2}&0\end{bmatrix}T^{-1}(\boldsymbol{\vartheta})-\begin{bmatrix}0\\ d\end{bmatrix}\right\}-\frac{\partial T}{\partial\boldsymbol{\zeta}}\begin{bmatrix}\displaystyle\frac{1}{|\zeta_{1}|}\\ 0\end{bmatrix}\rho(\tilde{u}-\hat{u}),$ (62)
$\exists V_{ST}=\boldsymbol{\vartheta}^{T}P_{ST}\,\boldsymbol{\vartheta}$ such that $V_{2}\leq-c||\boldsymbol{\vartheta}||\triangleq-W_{ST}(||\boldsymbol{\vartheta}||)$.

### v4

@@
-Dynamic Input Mapping Inversion for Algebraic Loop-Free Control in Hydraulic Actuators
+Dynamic Input Mapping Inversion to Eliminate Algebraic Loops in Hydraulic Actuator Control
 Abstract
 The application of nonlinear control schemes to electro-hydraulic actuators often requires several alterations in the design of the controllers during their implementation.
-This is to overcome the challenges that frequently arise from the inherent complexity of such control algorithms owning to model nonlinearities.
-Moreover, advanced control solutions for this type of systems often introduce input algebraic loops and chatter, which considerably degrade the tracking performance.
-This study presents a nonlinear control architecture for hydraulic actuators that comprises low-complexity modules, based on well-established designs that facilitate robust high performance in tracking without introducing the aforementioned limitations.
-Specifically, the proposed solution consists of two variants of a position controller for the hydraulic cylinder and a dynamic input-mapping inversion module to avoid algebraic loops in the control input.
-The stability of the closed-loop system is analysed using arguments from Lyapunov theory for cascaded non-autonomous nonlinear systems.
-The effectiveness of the proposed solution is evaluated on a high-fidelity simulator of a wind turbine pitch system.
-Appropriate quantitative metrics are finally defined to evaluate the closed-loop system performance in comparison to state-of-the-art nonlinear design.
-Index Terms:
-electro-hydraulic actuator, nonlinear control, parameter estimation, mapping inversion
+This is to overcome challenges that frequently arise in such control algorithms owing to model nonlinearities.
+Moreover, advanced control solutions for this type of systems often introduce input algebraic loops that pose significant design and tuning difficulties.
+Conventional methods to avoid such loops introduce chatter, which considerably degrade tracking performance and has oil degradation and wear as side effects.
+This study presents a nonlinear control architecture for hydraulic actuators that comprises low-complexity modules that facilitate robust high performance in tracking and avoids the drawbacks of chatter.
+The salient feature is a dynamic input-mapping inversion module that avoids algebraic loops in the control input and is followed by dedicated position control.
+The stability of the closed-loop system is analyzed using arguments from Lyapunov theory for cascaded non-autonomous nonlinear systems.
+The effectiveness of the proposed solution is evaluated on a high-fidelity simulator of a wind turbine pitch system, and validated on a full-scale laboratory setup that includes a hydraulic pitch system and blade bearing.
+Appropriate quantitative metrics are used to evaluate the closed-loop system performance in comparison to a state-of-the-art nonlinear design.
 I Introduction
 Electro-hydrualic actuators (EHA) are commonly used in industrial applications because of their high power density and resilience to harsh environment.
-Typical applications include aviation, manufacturing, marine/offshore engineering, robotics and energy systems [1, 2, 3].
+Typical applications include aviation, manufacturing, marine/offshore engineering, robotics and energy systems [fossenMarineControlSystems2002, walgernReliabilityElectricalHydraulic2023, dallabonaFaultDiagnosisPrognosis2025].
 Although Proportional-Integral-Differential (PID) controllers constitute the industrial standard in these application domains, they often fall short in maintaining high performance in the presence of strong nonlinearities and uncertainties that characterize hydraulic actuator systems.
 This has motivated the exploration of nonlinear control alternatives that can improve performance and enhance closed-loop robustness over the entire envelope of operation.
-The complexity, however, of system dynamics often hinders effective application of nonlinear model-based solutions such as feedback linearization (FL) or sliding mode control (SMC) because several limiting assumptions and modelling simplifications need be imposed to apply these methods.
+The complexity, however, of system dynamics often hinders effective application of nonlinear model-based solutions because several limiting assumptions and modeling simplifications need be imposed to apply these methods.
 A common heavy simplification concerns the valve dynamics, where two main issues can be identified.
-A first limitation is encountered when a system model cannot be written as an affine form $\dot{x}=f(x)+g(x)u$, as the vector field of a valve depends non-linearly on the input.
+A first limitation is encountered when a system model cannot be written in the affine form $\dot{x}=f(x)+g(x)u$, as the vector field depends non-linearly on the input.
 A common approach to address this issue is to use the input value at the previous time instant, under the assumption that the input $u$ is slowly varying.
-This removes the nonlinearity and avoids the introduction of an algebraic loop, at the cost of increased chatter in the control signal [4].
-The second design challenge arises with inclusion of actuator dynamics, which increase the relative degree of the system, thus making the design much more complicated, as will be discussed later.
+This removes the nonlinearity and avoids the introduction of an algebraic loop, at the cost of increased chatter in the control signal [ayalewCascadeTuningNonlinear2006].
+The second design challenge arises with inclusion of actuator dynamics, which increases the relative degree of the system, thus making the design much more complicated, as will be discussed later.
 Neglecting the valve dynamics during the design stage can simplify implementation but may potentially result in a quite demanding control action going beyond obtainable valve response.
@@
 Furthermore, micro-movements in EHAs are responsible for local heating of hydraulic oil, which degrades oil quality.
-This leads to increased friction, and eventually to wear and potential damage of a valve [5].
+This leads to increased friction, and eventually to wear and potential damage of the valve [dallabonaFrictionEstimationCondition2024a].
 It is therefore essential to establish control designs that combine the performance and robustness features of nonlinear solutions, but avoid excessive actuation demand.
-Several studies in the existing literature have reported solutions that adopt nonlinear strategies for controlling hydraulic systems.
-A cascaded feedback linearization controller with a sliding mode observer was designed in [6].
-Sliding mode was explored in a standard fashion in [7], in combination with adaptive control in [8], with an integral sliding surface in [9] and later in [10] to eliminate the velocity state in the control action.
+Several studies in existing literature have reported solutions that adopt nonlinear strategies for controlling hydraulic systems.
+A cascaded feedback linearization controller with a sliding mode observer was designed in [kochObserverbasedSlidingMode2016].
+Sliding mode was explored in a standard fashion in [thomasPracticalProfileTracking2022], [guanAdaptiveSlidingMode2008a], for high order sliding modes in [estradaHydraulicActuatorControl2025] and in [estradaSupertwistingBasedSliding2024] to eliminate the velocity state in the control action.
+The authors in [yamamotoForceControllerValvemanipulated2025] and [yamamotoPositionControllerHydraulic2025] combined sliding mode with a quasi-static EHA model using a set-valued formalism to capture model nonlinearities.
+A fixed-time Event-Triggered Control strategy has been tested in [wangDisturbanceObserverBasedFixedTime2025].
 Adaptive solutions have been applied by means of backstepping, in e.g.
-[11], [12], [13], and [14].
-The authors in [15] used a coordinate transformation and tuning functions to obtain a backstepping controller that included the valve dynamics.
-The high relative degree led to a complex solution.
-Except for the latter, all the other solution adopted the strategy presented in [4] to cancel the algebraic loop, and do not account for the valve dynamics in the design of the controller.
-The aim of this paper is to leverage the results on online estimation of nonlinearly parametrised perturbations [16, 17] as a dynamical mapping inversion (MI) algorithm to derive the control input from a virtual input designed via classical nonlinear methods.
+[chengguanNonlinearAdaptiveRobust2008], [chouxCascadeControllerIncluding2012], [bakhshandeProportionalIntegralObserverBasedBacksteppingApproach2018], [manganasBacksteppingBasedController2023], and neural networks in [guoNeuralAdaptiveBackstepping2019], [yangNeuralAdaptiveDynamic2023], [phanAdaptiveNeuralObserverbased2025].
+The authors in [chouxAdaptiveBacksteppingControl2010a] used a coordinate transformation and tuning functions to obtain a backstepping controller that included the valve dynamics.
+The high relative degree led to a complex solution that would also be affected by valve degradation over time.
+Except for said article, all other published solutions adopted the strategy presented in [ayalewCascadeTuningNonlinear2006] to cancel the algebraic loop, and did not account for the input valve dynamics in the design of the controller.
+This paper uses results from online estimation of nonlinearly parametrized perturbations [gripParameterEstimationCompensation2010b, papageorgiouRobustBacklashEstimation2019] to obtain a dynamical mapping inversion (MI) algorithm and derive the control input from a virtual input designed via classical nonlinear methods, without relying on the simplifying assumptions commonly adopted in the literature to implement nonlinear designs.
 The paper shows how a modular and straightforward design is achieved, and how the MI can reconstruct the control signal without introducing algebraic loops.
 Moreover, it is shown how the suggested approach facilitates a systematic commissioning feature, namely, how it can accommodate valve specification parameters such that only frequencies that can actually be delivered by the valve are allowed in the control input.
-In this way, the performance can be improved and chatter in the input be significantly reduced.
-This article contributes by
+In this way, the performance can be improved and chatter in the input can be significantly reduced.
+The contributions of this article are:
   1. 1.
-    presenting a modular architecture for nonlinear control of hydraulic actuators without introducing algebraic loops or limiting assumptions on the input signal,
+    A dynamical input mapping inversion algorithm in a modular architecture for control of hydraulic actuators that avoids algebraic loops and significantly reduces EHA micro-movements and associated wear.
   2. 2.
-    proposing a scheme for reconstructing the control signal by leveraging a nonlinear parameter estimation algorithm for dynamically inverting the mapping between control input and virtual input,
+    A systematic method to tune the dynamical mapping inversion component according to control valve dynamics.
   3. 3.
-    introducing a systematic method to tune the dynamical mapping inversion component according to the actuator parameters,
-  4. 4.
-    providing rigorous stability and boundedness analyses of the closed-loop dynamics under different control laws,
-  5. 5.
-    demonstrating the effectiveness of the proposed solution against conventional nonlinear control methods for industrial EHA systems in a real case scenario, using a high-fidelity simulator.
+    Rigorous analysis of stability/boundedness and validation on a wind-turbine EHA system, using both a high-fidelity simulator and a full-scale physical test rig.
 The remainder of the paper is organized as follows.
@@
 Sec.
-III designs controllers following commonly accepted standards, and Sec.
+III designs a well-established baseline controller following commonly accepted standards, and Sec.
 IV proposes a new algorithm and presents features and benefits in contrast to the standard existing approaches.
@@
 Sec.
-VI presents a case study of a wind turbine pitch control, and results from high fidelity simulation are shown in Sec.
+VI presents a case study of a wind turbine pitch control, and results from validation on a full-scale electro-hydraulic test rig are shown in Sec.
 VII.
-Finally, conclusions and future directions are staged in Sec.
+Finally, conclusions and future directions are offered in Sec.
 VIII.
 II Problem definition
-An electro-hydraulic actuator is composed of an hydraulic cylinder driven by a directional control valve, as shown in Fig.
+An electro-hydraulic actuator is composed of an hydraulic cylinder driven by a directional control valve (DCV), as shown in Fig.
 1.
-The pressures $p_{p}$ and $p_{r}$ in the piston-side and rod-side chambers of the cylinder are varied according to the valve spool position $x_{v}$.
+The pressures $p_{p}$ and $p_{r}$ in the piston-side and rod-side chambers of the cylinder are varied according to the valve’s spool position $x_{v}$.
 This allows for control of both position and velocity ($x_{c}$ and $v_{c}$) of the the cylinder’s piston.
-The model for a generic electro-hydraulic actuator is given by [9]:
+The model for a generic electro-hydraulic actuator is given by the dynamic model 1, see e.g.
+[komstaIntegralSlidingMode2013, dallabonaFaultDiagnosisPrognosis2025, pedersenInvestigationLoadReduction2016]:
 $M\dot{v}_{c}=A_{p}p_{p}-A_{r}p_{r}-F_{Fr}-F_{ext}$ (1)
@@
 $\dot{x_{v}}=v_{v},$
-Figure 1: Generic electro-hydraulic actuator configuration: above, the linear cylinder with piston and rod; below, the flow control valve, oil pump and drain.
+Figure 1: Generic electro-hydraulic actuator system: above, the linear cylinder with piston and rod; below, the flow control valve, oil pump and drain.
 where $M$ is the piston mass, $A_{p}$ and $A_{r}$ the cylinder cross sections in the two chambers, $F_{fr}$ and $F_{ext}$ the friction force and the external load respectively, and $B_{e,i}(p_{i})$ the effective bulk modulus of the oil in the $i$-th component.
-The control input $u_{x}$ is a position reference for the control valve’s spool.
 The terms
@@
 $\Delta_{pr}=(p_{r}-p_{t})\,H\left(x_{v}\right)+(p_{s}-p_{r})\,H\left(-x_{v}\right),$
-with $H(\cdot)$ being the Heaviside step function.
+where $H(\cdot)$ stands for the Heaviside step function, and $p_{s}$ and $p_{t}$ for system and tank pressures, respectively.
+The terms $\omega_{0}$ and $\xi$ denote the natural frequency and damping ratio of the input valve, which is modeled as a second-order system.
+This stems from DCVs typically being equipped with an internal closed loop.
+As a result, the control input $u_{x}$ is a normalized, thus dimensionless, position reference for the valve spool.
 The control objective in EHA systems is the tracking of a reference signal for the cylinder piston position $x_{c}$.
-By inspecting the cylinder position dynamics in Eq.
-1 it is easily seen that the system corresponding to the position error defined as
-$e\triangleq x_{c}-r,$ (5)
-has relative degree 5, i.e. $e$ needs to be differentiated five times in order to arrive to a controller canonical form
+By inspecting the cylinder position dynamics in (1) it is easily seen that the system corresponding to the position tracking error $e$,
+$e=x_{c}-r,$ (5)
+has relative degree $\alpha=5$, i.e. $e$ needs to be differentiated five times in order to arrive to a controller canonical form
 $e^{(\alpha)}=f(u),$ (6)
@@
 For the class of EHA systems this introduces a high degree of complexity due to the strong nonlinear couplings between the various pressure terms.
-A common approach adopted in the literature to alleviate this complexity is the introduction of a simplifying assumption, namely the neglection of the valve dynamics.
+A common approach adopted in the literature to alleviate this complexity is the introduction of a simplifying assumption, namely neglecting the valve dynamics.
 This implies that the position $x_{v}$ inside the terms $Q_{p}(x_{v})$ and $Q_{r}(x_{v})$ can instantaneously assume any desired value and is, therefore, considered the effective control input for the system, i.e.
-$u\triangleq x_{v}$, with $u\in[-1,1]$.
+$u\triangleq x_{v}$ and $Q_{p}(x_{v}),Q_{r}(x_{v})$ becoming $Q_{p}(u),Q_{r}(u)$.
 An additional frequent simplification made in many reported studies is considering one single constant value for the bulk modulus of the oil for both chambers, i.e., $B_{e,p}=B_{e,r}=B>0$.
-The resulting model, which is most commonly used in the industry for control design, is [18]:
+The resulting model, which is most commonly used in the industry for control design, is [bakhshandeRobustControlApproach2017]:
 $M\dot{v}_{c}=A_{p}p_{p}-A_{r}p_{r}-F_{Fr}-F_{ext}$ (7)
@@
 III Baseline nonlinear control design
-Considering the model presented in Eq. 7 and the position error $e$ defined in Eq. 5, the dynamics constitutes a relative degree 3 system and reads:
+The control architecture for the baseline solution is presented in Fig. 2, and represent the standard configuration for nonlinear control solutions for the system under analysis.
+Considering the model presented in (7) and the tracking error $e$ defined in (5), the dynamics constitutes a relative degree 3 system and reads,
 $e^{(3)}=$ $-\frac{\dot{F}_{fric}+\dot{F}_{ext}}{M}-\frac{A_{p}^{2}B}{MV_{p}}v_{c}-\frac{A_{r}^{2}B}{MV_{r}}v_{c}-r^{(3)}+$ (8)
-$\frac{A_{p}B}{MV_{p}}Q_{p}(u)+\frac{A_{r}B}{MV_{r}}Q_{r}(u)\triangleq\,\varpi_{FL}.$
+$\frac{A_{p}B}{MV_{p}}Q_{p}(u)+\frac{A_{r}B}{MV_{r}}Q_{r}(u)\triangleq\,\varpi.$
+where the term $\varpi=\varpi(t,\boldsymbol{x},u)$ is a function of time, of states $\boldsymbol{x}\triangleq\begin{bmatrix}v_{c}&x_{c}&F_{h}\end{bmatrix}^{T}$, and of the control input $u$.
 Intermediate derivatives are given by
 $\dot{e}$ $=v_{c}-\dot{r}$ (9)
-$\ddot{e}$ $=-\frac{F_{fric}+F_{ext}}{M}+\frac{A_{p}p_{p}-A_{r}p_{r}}{M}-\ddot{r},$
-where the virtual input $\varpi_{FL}=\varpi_{FL}(t,\boldsymbol{x},u)$ is a function of time, of states $\boldsymbol{x}\triangleq\begin{bmatrix}v_{c}&x_{c}&F_{h}\end{bmatrix}^{T}$, and of the control input $u$.
-Two basic control solutions are first presented to demonstrate both the effectiveness and the reasons behind the development of a new algorithm.
-Feedback linearization is presented as a baseline nonlinear control design, which is popular for adaptive nonlinear control solutions.
-A supertwisting sliding mode control design is shown as a second popular solution.
-III-A Feedback linearization (FL)
-A FL controller is obtained from writing the error derivative dynamics as follows,
-$\dot{\underline{\mathbf{e}}}=\underbrace{\begin{bmatrix}0&1&0\\ 0&0&1\\ 0&0&0\end{bmatrix}}_{\boldsymbol{A_{FL}}}\underbrace{\begin{bmatrix}e\\ \dot{e}\\ \ddot{e}\end{bmatrix}}_{\underline{\mathbf{e}}}+\underbrace{\begin{bmatrix}0\\ 0\\ 1\end{bmatrix}}_{\boldsymbol{B_{FL}}}\varpi_{FL}.$ (10)
-Selecting
-$\varpi_{FL}=-K_{FL}\underline{\mathbf{e}}$ (11)
-with $K_{FL}$ obtained with a standard linear design (eignvalue assignment, LQR, etc.), renders the closed-loop system globally exponentially stable (GES) with the rate of decay being defined by the gain $K_{FL}$.
-The related control input $u$ is isolated from Eq.
-11 taking also Eq.
-8, and Eq.
-9 into account.
-III-B Supertwisting SMC
-Sliding mode controllers aim at enforcing the trajectories $\boldsymbol{e}(t)\in\mathbb{R}^{3}$ of the error system onto a pre-designed manifold $\mathcal{S}=\{\boldsymbol{e}\in\mathbb{R}^{3}\,\big{|}\,s(\boldsymbol{e})=0\}$.
-The variable $s$ is referred to as the sliding variable and is designed to guarantee exponential stability of the error system, so long $s$ is regulated at the origin.
-In order to leverage the reduced chatter properties of the well-studied supertwisting sliding mode control designs [19], the dynamics of s should be brought into the form of Eq.
-6 with a single differentiation.
-Therefore, the sliding variable $s$ is defined as:
-$s=\ddot{e}+\lambda_{2}\dot{e}+\lambda_{1}e,$ (12)
-where $\lambda_{1},\lambda_{2}$ are selected such that the characteristic polynomial in Eq. 12 is Hurwitz. The dynamics of the sliding variable $s$ is:
-$\dot{s}=\underbrace{e^{(3)}+\lambda_{2}\ddot{e}+\lambda_{2}\dot{e}}_{\displaystyle\varpi_{ST}(t,\boldsymbol{x},u)},$ (13)
-where the right-hand side is of Eq.
-13 is again a function $\varpi_{ST}$ of time, the states and the control input $u$.
-Eq.
-12 implies global exponential stability for the origin of the piston position error when the system is in sliding mode, i.e.
-when $s=\dot{s}=0$.
-This can be achieved by selecting
-$\varpi_{ST}=-k_{1}\lfloor s\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s(\tau))d\tau,$ (14)
-where $k_{1},k_{2}>0$. Similar to the controller case in the previous subsection, the control input is derived by isolating the term $u$ in equation 14.
-III-C Control input
-The final step for designing the controllers is deriving the control input $u$ for which the closed-loop system dynamics is shaped, via the previously presented design methods.
-From Eq.
-8 it is noticed that the error dynamics depends on the terms $F_{ext}$ and $F_{fr}$, and their derivatives.
-Such terms are in general unknown and need be estimated to compute the control input $u$.
-III-C1 Disturbance observer
-Although dedicated tests can be run to identify the friction coefficients, it is convenient in this investigation to assume the friction term to be unknown.
-This allows for lumping it together with the external load in a single term $F_{d}=F_{ext}+F_{fr}$, which can then be estimated with a conventional linear observer.
-Such an observer can be designed based on the measured pressures as follows,
-$\dot{\hat{\boldsymbol{z}}}=\underbrace{\begin{bmatrix}0&-\frac{1}{M}\\ 0&0\end{bmatrix}}_{\boldsymbol{A}}\hat{\boldsymbol{z}}+\begin{bmatrix}\displaystyle\frac{A_{p}p_{p}-A_{r}p_{r}}{M}\\ 0\end{bmatrix}+\underbrace{\begin{bmatrix}L_{1}\\ L_{2}\end{bmatrix}}_{\boldsymbol{L}}\underbrace{\begin{bmatrix}1&0\end{bmatrix}}_{\boldsymbol{C}}\tilde{\boldsymbol{z}},$ (15)
-where $\hat{\boldsymbol{z}}=\begin{bmatrix}\hat{v_{c}}&\hat{F}_{d}\end{bmatrix}^{T}$, $\tilde{\boldsymbol{z}}\triangleq\boldsymbol{z}-\hat{\boldsymbol{z}}$ and $\boldsymbol{L}$ chosen such that $\boldsymbol{A}-\boldsymbol{LC}$ is Hurwitz, ensuring that the disturbance term $F_{d}$ can be reconstructed with a bounded error [20].
-Furthermore, since $\boldsymbol{z}\in\mathcal{L}_{\infty}$ and $\boldsymbol{A}-\boldsymbol{LC}$ is Hurwitz by design, it holds that
-$\|\boldsymbol{\dot{\tilde{z}}}\|=\|(\boldsymbol{A}-\boldsymbol{LC})\,\boldsymbol{\tilde{z}}\|\leq-\lambda_{min}(\boldsymbol{A}-\boldsymbol{LC})\,\|\boldsymbol{\tilde{z}}\|_{\infty}\triangleq L_{z},$ (16)
-where $L_{z}$ is a positive real number. Thus, $\dot{F}_{d}$ is also reconstructed with a bounded error. Global asymptotic stability of the entire closed-loop system will be proved subsequently.
-III-C2 Control input derivation
-Directly isolating $u$ in Eq.
-11 and Eq.
-13 is not possible: while it may be possible to invert the $K(u)$ function in Eq.
-3, the control input also appears inside the Heaviside functions, generating an algebraic loop where the derived input depends on the input itself.
-The standard way to solve the problem is the one described in [4]:
-$Q_{p}(t,u)=$ $K(u)\left[\sqrt{p_{s}-p_{p}}\,H\left(u\right)+\sqrt{p_{p}-p_{t}}\,H\left(-u\right)\right]$ (17)
-$\approx$ $Ku\left[\sqrt{p_{s}-p_{p}}\,H\left(u_{-1}\right)+\sqrt{p_{p}-p_{t}}\,H\left(-u_{-1}\right)\right]$
+$\ddot{e}$ $=-\frac{F_{fr}+F_{ext}}{M}+\frac{A_{p}p_{p}-A_{r}p_{r}}{M}-\ddot{r},$
+Figure 2: Block diagram of the baseline algorithm that does not consider DCV dynamics.
+Feedback linearization (FL) is chosen as the baseline nonlinear control design to showcase the benefits of the new developed algorithm, given its widespread adoption in adaptive control schemes.
+Other solutions like sliding mode control, which often rely on an equivalent control law for satisfactory performance[dallabonaModelfreeSlidingMode2024], could similarly be used.
+A feedback linearization (FL) control is obtained from writing the error derivative dynamics as follows,
+$\dot{\underline{\mathbf{e}}}=\underbrace{\begin{bmatrix}0&1&0\\ 0&0&1\\ 0&0&0\end{bmatrix}}_{\boldsymbol{A_{FL}}}\underbrace{\begin{bmatrix}e\\ \dot{e}\\ \ddot{e}\end{bmatrix}}_{\underline{\mathbf{e}}}+\underbrace{\begin{bmatrix}0\\ 0\\ 1\end{bmatrix}}_{\boldsymbol{B_{FL}}}\varpi.$ (10)
+Selecting $\varpi$ as
+$\varpi=-K_{FL}\underline{\mathbf{e}},$ (11)
+with $K_{FL}$ obtained by a standard linear design (eigenvalue assignment, LQR, or other), renders the closed-loop system globally exponentially stable (GES) with the rate of decay being defined by the gain $K_{FL}$.
+The related control input $u$ is isolated from (11) taking also (8), and (9) into account.
+From the latter it is noticed that the error dynamics depends on the terms $F_{ext}$ and $F_{fr}$, and their derivatives.
+Such terms are in general unknown and need be estimated to later compute the control input $u$.
+Although dedicated offline tests could identify the friction coefficients as they are during commissioning, the friction will change over time and it is both convenient and necessary to assume friction to be unknown.
+This allows for lumping friction together with external load as a single term $F_{d}=F_{ext}+F_{fr}$, which can be estimated by a conventional observer, using pressure measurements.
+One possible solution to this estimation problem was presented in [dallabonaModelfreeSlidingMode2024].
+III-A Standard control input derivation
+Directly isolating $u$ in (11) is not possible: while it may be possible to invert the $K(u)$ function in (3), the control input also appears inside the Heaviside functions, generating an algebraic loop where the derived input depends on the input itself.
+The standard way to solve this problem was described by [ayalewCascadeTuningNonlinear2006],:
+$Q_{p}(t,u)=$ $K(u)\left[\sqrt{(p_{s}-p_{p})\,H(u)}+\sqrt{(p_{p}-p_{t})\,H(-u)}\right]$ (12)
+$\approx$ $Ku\left[\sqrt{(p_{s}-p_{p})\,H(u_{-1})}+\sqrt{(p_{p}-p_{t})\,H(-u_{-1})}\right]$
 $=$ $K_{Qp}\,u,$
-$Q_{r}(t,u)=$ $K(u)\left[\sqrt{p_{r}-p_{t}}\,H\left(u\right)+\sqrt{p_{s}-p_{r}}\,H\left(-u\right)\right]$ (18)
-$\approx$ $Ku\left[\sqrt{p_{r}-p_{t}}\,H\left(u_{-1}\right)+\sqrt{p_{s}-p_{r}}\,H\left(-u_{-1}\right)\right]$
+$Q_{r}(t,u)=$ $K(u)\left[\sqrt{(p_{r}-p_{t})\,H(u)}+\sqrt{(p_{s}-p_{r})\,H(-u)}\right]$ (13)
+$\approx$ $Ku\left[\sqrt{(p_{r}-p_{t})\,H(u_{-1})}+\sqrt{(p_{s}-p_{r})\,H(-u_{-1})}\right]$
 $=$ $K_{Qr}\,u,$
@@
 $g=\frac{A_{p}B}{MV_{p}}K_{Qp}+\frac{A_{r}B}{MV_{r}}K_{Qr},$ (14)
-the standard feedback linearization controller is then,
-$u_{FL}=\frac{1}{g}\left[\frac{\dot{\hat{F}}_{d}}{M}+\frac{A_{p}^{2}B}{MV_{p}}v_{c}+\frac{A_{r}^{2}B}{MV_{r}}v_{c}+r^{(3)}-K_{FL}\underline{\mathbf{e}}\right],$ (20)
-and the standard super-twisting sliding mode controller is,
-$u_{ST}=\frac{1}{g}$ $\left[\frac{\dot{\hat{F}}_{d}}{M}+\frac{A_{p}^{2}B}{MV_{p}}v_{c}+\frac{A_{r}^{2}B}{MV_{r}}v_{c}-\lambda_{2}\ddot{e}-\lambda_{1}\dot{e}\right.$ (21)
-$\,\,\left.+r^{(3)}-k_{1}\lfloor s\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s)d\tau\right].$
-The $K(u)$ linearization is then reflected into uncertainties in the input term $g$ and in the terms $K_{Qp}$ and $K_{Qr}$.
+a standard feedback linearizing control is obtained by,
+$u=\frac{1}{g}\left[\frac{\dot{\hat{F}}_{d}}{M}+\frac{A_{p}^{2}B}{MV_{p}}v_{c}+\frac{A_{r}^{2}B}{MV_{r}}v_{c}+r^{(3)}-K_{FL}\underline{\mathbf{e}}\right].$ (15)
+The $K(u)$ linearization is then reflected into uncertainties in the input term $g$ and in the terms $K_{Qp}$ and $K_{Qr}$ from (12) and (13).
 IV Cascaded control design
-The delay operator introduced to avoid the formation of algebraic loops simplified the controller implementation with the drawback of increasing unwanted chatter on the control signal, with the consequences on oil quality described above.
-The solution proposed in this article for avoiding this, is to leverage a virtual input for the controller design and dynamically deriving a real input that produces the virtual one required by the controller.
-The block diagram is shown in Fig.
-2.
-Figure 2: Block diagram of the proposed algorithm: Cascaded control with dynamic input mapping inversion.
-By defining a virtual input as
+The delay operator introduced to avoid the formation of algebraic loops simplified the controller implementation but have the drawback of increasing unwanted chatter on the control signal [ayalewCascadeTuningNonlinear2006],with associated consequences on oil quality degradation that were mentioned above.
+The solution proposed in this article, to avoid this, is to leverage a virtual input for the controller design and dynamically deriving a real input that produces the virtual one required by the controller.
+The block diagram for the proposed algorithm is shown in Fig.
+3.
+Defining a virtual input
 $v\triangleq\frac{A_{p}B}{MV_{p}}Q_{p}(u)+\frac{A_{r}B}{MV_{r}}Q_{r}(u),$ (16)
-Eq. 8 can be rewritten as
+(8) can be rewritten as
 $e^{(3)}=-\frac{\dot{F}_{fric}+\dot{F}_{ext}}{M}-\frac{A_{p}^{2}B}{MV_{p}}v_{c}-\frac{A_{r}^{2}B}{MV_{r}}v_{c}-r^{(3)}+v,$ (17)
-where the virtual input $v$ is a function of time and the control input $u$, and is now present as a linear term.
-The feedback linearizing and supertwisting control designs are now defined in terms of the associated virtual inputs as, respectively,
-$v_{FL}=\frac{\dot{\hat{F}}_{d}}{M}+\frac{A_{p}^{2}B}{MV_{p}}v_{c}+\frac{A_{r}^{2}B}{MV_{r}}v_{c}+r^{(3)}-K_{FL}\underline{\mathbf{e}},$ (24)
-and
-$v_{ST}=$ $\frac{\dot{\hat{F}}_{d}}{M}+\frac{A_{p}^{2}B}{MV_{p}}v_{c}+\frac{A_{r}^{2}B}{MV_{r}}v_{c}-\lambda_{2}\ddot{e}-\lambda_{1}\dot{e}$ (25)
-$+r^{(3)}-k_{1}\lfloor s\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s(\tau))\,d\tau.$
+where $v$ is a function of time and the control input $u$, and is now present as a linear term. A feedback linearizing controller is now defined in terms of an associated virtual input as
+$v=\frac{\dot{\hat{F}}_{d}}{M}+\frac{A_{p}^{2}B}{MV_{p}}v_{c}+\frac{A_{r}^{2}B}{MV_{r}}v_{c}+r^{(3)}-K_{FL}\underline{\mathbf{e}}.$ (18)
 IV-A Dynamic mapping inversion algorithm
-The last part of the proposed method constitutes the derivation of the control input $u$ by solving Eq.
-22 for $u$ in each case.
-A dynamic nonlinear mapping inversion scheme is adopted from [16] for this purpose, allowing for re-formulating of the input calculation into a parameter estimation problem in $u$ for a nonlinearly parametrized regressor.
+The last part of the proposed method constitutes the derivation of the control input by solving (16) for $u$.
+A dynamic nonlinear mapping inversion scheme is adopted from [gripParameterEstimationCompensation2010b] for this purpose, allowing for re-formulating of the input calculation into a parameter estimation problem in $u$ for a nonlinearly parametrized regressor.
 Specifically, let $\rho:[0,\infty)\times\mathbb{R}^{3}\times[-1,1]\rightarrow\mathbb{R}$ with
 $\rho(t,\boldsymbol{x},u)=\frac{A_{p}B}{MV_{p}}Q_{p}(t,\boldsymbol{x},u)+\frac{A_{r}B}{MV_{r}}Q_{r}(t,\boldsymbol{x},u)$ (19)
-be a mapping from any given actual control input $u$ to the virtual input that was earlier defined in Eq.
-22.
+Figure 3: Block diagram of the proposed algorithm: Cascaded control with dynamic input mapping inversion. The input inversion is explained below.
+be a mapping from any given actual control input $u$ to the virtual input that was defined in (16).
 Assume that for every value $v$ at each time instant there exists an appropriate value of the control input $u^{*}$ such that $v=\rho(u^{*})$.
-It has been shown [16, 17] that the estimation law
+It has been shown [gripParameterEstimationCompensation2010b, papageorgiouRobustBacklashEstimation2019] that the estimation law
 $\dot{u}=\gamma\,\mu(t,\boldsymbol{x},u)\,\Big(v-\rho(t,\boldsymbol{x},u)\Big),$ (20)
@@
 There exist positive constants $T,\alpha_{0}>0$ such that
-[…]
+$\int_{t}^{t+T}\sigma(\tau,\boldsymbol{x}(\tau))d\tau\geq\alpha_{0}\boldsymbol{I}\;,$ (23)
 where $\boldsymbol{I}$ is the identity matrix.
+Figure 4: Manufacturer’s lookup table of valve characteristic and approximation with a polynomial function.
+The final control action is then given by
+$u(t)=\gamma\int_{0}^{t}\mu\big(\boldsymbol{x}(\tau),u(\tau)\big)\,\Big(v(\tau)-\rho\big(\boldsymbol{x}(\tau),u(\tau)\big)\Big)d\tau,$ (24)
+where $v$ is the one defined in (18).
 Remark 1.
-If the mapping $\rho$ is not injective, then under Conditions 1-3, the estimation law in Eq.
-27 will ensure that $u$ converges to an appropriate value such that $\rho(t,u)$ converges to $v$ exponentially fast.
-Figure 3: Lookup table approximation with polynomial functions.
+If the mapping $\rho$ is not injective, then under Conditions 1-3, the estimation law in (20) will ensure that $u$ converges to an appropriate value such that $\rho(t,u)$ converges to $v$ exponentially fast.
 A typical choice for $\mu$ is the gradient of $\rho$ with respect to $u$.
-However, computation of $\mu$ requires differentiation of the nonlinear function $K(u)$ present in Eq.
-3 and, since the latter is usually available as a lookup table, it is approximated by an appropriate polynomial function of $u$.
+However, computation of $\mu$ requires differentiation of the nonlinear function $K(u)$ present in (3) and, since the latter is usually available as a lookup table, it is approximated by an appropriate polynomial function of $u$.
 Fig.
-4 illustrates an example of different polynomial approximations for $K(u)$, i.e.
-$K(u)=\sum_{i=0}^{n}\alpha_{i}\,u^{i},$ (31)
-with a fifth-order polynomial yielding the best fit.
-This particular approximation also ensures that $\frac{\partial K}{\partial u}>~{}0,\;\forall u\in[-1,1]$ as shown in Fig.
-4, which will be useful later in proving the validity of the convergence conditions.
+5 illustrates a ninth-order polynomial approximation for $K(u)$, i.e.
+$K(u)=\sum_{i=0}^{9}\alpha_{i}\,u^{i},$ (25)
+for both an ideal actuator and the deteriorated (denoted real in Fig 5) actuator used later in the experimental validation.
+Such approximation also ensures that for the case of the ideal valve $\frac{\partial K}{\partial u}>~0,\;\forall u\in[-1,1]$, as shown in Fig.
+5.
+This condition will be useful later in proving the validity of the convergence conditions.
 The function of $\mu$ can now be selected as
@@
 $\geq 2\left(\inf\limits_{u}\mu(t,\boldsymbol{x},u)\right)^{2}\triangleq\sigma(t,\boldsymbol{x})>0\;.$
-Moreover, since $\frac{\partial\rho}{\partial u}$ is bounded, there exists a Lipschitz constant $L_{u}>0$ such that
+Moreover, since $\left|\displaystyle\frac{\partial\rho}{\partial u}\right|$ is bounded, there exists a Lipschitz constant $L_{u}>0$ such that
 $\left|\rho(t,\boldsymbol{x},u^{*})-\rho(t,\boldsymbol{x},u)\right|<L_{u}|\tilde{u}|\leq\frac{L_{u}|\tilde{u}|\sqrt{\sigma(t,\boldsymbol{x})}}{\sqrt{\inf\limits_{(t,\boldsymbol{x})}\sigma(t,\boldsymbol{x})}}$ (28)
@@
 Hence, it would only be reflected in additional spikes in case a discontinuity is present during switching.
-Figure 4: Verification of derivative for algorithm convergence condition.
+Figure 5: Flow coefficient derivative for algorithm convergence condition and valve degradation effect.
+Remark 3.
+In the case of a non-ideal valve, e.g.
+due to wear, the term $\frac{\partial K}{\partial u}$ may take non-positive values as illustrated in Fig.
+5, which may lead to violation of Condition 1.
+This will, in turn, result to a bounded input estimation error.
+An alternative yet more conservative choice is to select $\mu$ to be a positive constant.
 V Stability Analysis
 The stability of the closed-loop system is analyzed by including the nonlinear controller, the dynamical mapping inversion, and the disturbance estimator.
-By writing it in the form of a feedback interconnected system
+Let $\delta_{\underline{\mathbf{e}}}$ be the error associated with the calculation of $\ddot{e}$.
+The dynamics of the closed-loop tracking error $\underline{\mathbf{e}}$ is written as:
 […]
-and by fulfilling the following conditions:
-Condition 4.
-Unperturbed $\Sigma_{1}$ $(\tilde{u}=0)$ is Uniformly GES
-Condition 5.
-Solutions of $\Sigma_{2}$ are Uniformly Ultimately Bounded (UUB)
-Condition 6.
-$\forall(t,\boldsymbol{\zeta})\in[0,\infty)\times\mathbf{R}^{2}$
-$\frac{\partial V(t,\boldsymbol{\zeta})}{\partial\boldsymbol{\zeta}}h(t,\tilde{u})=o(W(\boldsymbol{\zeta})),$ (36)
-the overall closed-loop system is globally UUB according to Theorem 1 from [21].
-Let $v^{*}=\rho(u^{*})$ be the required virtual controller and $u^{*}$ the related control input, one can define
-$\tilde{v}\triangleq v^{*}-v=\rho(u^{*})-\rho(u),$ (37)
-for some arbitrary $u$ and $v$, with $\tilde{u}=u^{*}-u$ having been defined earlier.
-The two closed-loop systems obtained by closing the loop with the FL controller and the ST-SMC controller are, respectively:
-$\Sigma_{1,FL}:\dot{\underline{\mathbf{e}}}=(\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}}K_{fl})\,\underline{\mathbf{e}}+\boldsymbol{\xi_{F_{d}}}+\boldsymbol{h_{FL}}(t,\tilde{u})$ (38)
-$\Sigma_{2,FL}:\dot{\tilde{u}}=\dot{u}^{*}-\gamma\,\mu(x,u^{*},\tilde{u})\,\Big{(}-K_{FL}\underline{\mathbf{e}}-\rho(u^{*}-\tilde{u})\Big{)}$
+[…]
+$=$ $(\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}K_{fl}})\,\underline{\mathbf{e}}+\boldsymbol{B_{FL}}\left(\boldsymbol{K_{FL}}\begin{bmatrix}0\\ -\frac{\tilde{F}_{d}}{M}\end{bmatrix}-\frac{\dot{\tilde{F}}_{d}}{M}\right)$
+[…]
+[…]
 with
-$\boldsymbol{\xi}_{F_{d}}=\boldsymbol{B_{FL}}\left(-\frac{\dot{\tilde{F}}_{d}}{M}+K_{FL}\begin{bmatrix}0\\ -\displaystyle\frac{\tilde{F}_{d}}{M}\end{bmatrix}\right)$ (39)
-$\boldsymbol{h_{FL}}(t,\tilde{u})=-\boldsymbol{B_{FL}}\,\rho(\tilde{u}),$
-and
-$\Sigma_{1,ST}:\dot{\boldsymbol{\vartheta}}=f_{\boldsymbol{\vartheta}}(t,\boldsymbol{\vartheta})+\boldsymbol{h_{ST}}(t,\tilde{u})$ (40)
-$\Sigma_{2,ST}:\dot{\tilde{u}}=\dot{u}^{*}-\gamma\,\mu(x,u^{*},\tilde{u})\,\Big{(}v^{*}(t,\boldsymbol{\vartheta})-\rho(u^{*}-\tilde{u})\Big{)}$
-with
-$f_{\boldsymbol{\vartheta}}(t,\boldsymbol{\vartheta})$ $=\frac{\partial T}{\partial\boldsymbol{\zeta}}\left\{\frac{1}{|\zeta_{1}|}\begin{bmatrix}-\frac{k_{1}}{2}&1\\ -k_{2}&0\end{bmatrix}T^{-1}(\boldsymbol{\vartheta})-\begin{bmatrix}0\\ d\end{bmatrix}\right\}$ (41)
-$\boldsymbol{h_{ST}}(t,\tilde{u})$ $=-\frac{\partial T}{\partial\boldsymbol{\zeta}}\begin{bmatrix}\displaystyle\frac{1}{|\zeta_{1}|}\\ 0\end{bmatrix}\rho(\tilde{u}-\hat{u})$
-according to [22], where $v^{*}(t,\boldsymbol{\zeta})$ was defined in Eq.
-25 and $\boldsymbol{\zeta}$ corresponds to a coordinates transformation of the closed-loop dynamics.
-The detailed derivations are presented in Appendix.
-Condition 4 coincides with the proof of stability of the controllers in their standard configuration, and is ensured by choosing the Lyapunov functions as:
-$V_{FL}$ $=\underline{\mathbf{e}}^{T}P_{FL}\,\underline{\mathbf{e}}$ (42)
-$V_{ST}$ $=\boldsymbol{\vartheta}^{T}P_{ST}\,\boldsymbol{\vartheta}.$
-Condition 5 is satisfied according to the result in [23], where it was proven that if $|\dot{u}^{*}|<\infty\,\forall t>0$, then $\tilde{u}(t)$ is UUB for all time.
-Finally, Condition 6 is verified as, $\Sigma_{1,FL}$:
-$\left\|\frac{\partial V_{FL}}{\partial\underline{\mathbf{e}}}\boldsymbol{h_{FL}}(t,\tilde{u})\right\|$ $\leq\left\|2P_{FL}\underline{\mathbf{e}}\right\|\,\left\|B_{FL}\right\|\,\left|\rho(u^{*})-\rho(u^{*}-\tilde{u})\right|$ (43)
-$\leq\left\|2P_{FL}\underline{\mathbf{e}}\right\|\,\left\|B_{FL}\right\|\,L_{p}\,\left\|\tilde{u}\right\|$
-$\leq\underbrace{\left\|2P_{FL}\right\|\,\left\|B_{FL}\right\|\,L_{p}\,\left\|\tilde{u}\right\|_{\infty}}_{\eta}\,\left\|\underline{\mathbf{e}}\right\|,$
-and for $\Sigma_{1,ST}$,
-$\left\|\frac{\partial V_{ST}}{\partial\boldsymbol{\vartheta}}\boldsymbol{h_{ST}}(t,\tilde{u})\right\|$ $\leq\left\|2P_{ST}\boldsymbol{\vartheta}\right\|\frac{L_{\rho}}{\min_{t}(|\zeta_{1}(t)|)}\|\tilde{u}\|_{\infty}\left\|\frac{\partial\tau}{\partial\boldsymbol{\xi}}\right\|_{\infty}$ (44)
-$\leq\underbrace{\left\|2P_{ST}\right\|\frac{L_{\rho}}{\min_{t}(|\zeta_{1}(t)|)}\|\tilde{u}\|_{\infty}\left\|\frac{\partial\tau}{\partial\boldsymbol{\xi}}\right\|_{\infty}}_{\eta^{\prime}}\left\|\boldsymbol{\vartheta}\right\|.$
+[…]
+[…]
+Moreover, let $v^{*}=\rho(u^{*})$ be the required virtual controller and $u^{*}$ the related control input. Then, one can define
+[…]
+for some arbitrary $u$ and $v$. From (20) the dynamics of the input estimation error $\tilde{u}$ reads:
+[…]
+The entire closed-loop system can therefore be written as
+$\underbrace{\begin{bmatrix}\dot{\underline{\mathbf{e}}}\\ \dot{\tilde{u}}\end{bmatrix}}_{\boldsymbol{\dot{\chi}}}=\underbrace{\begin{bmatrix}(\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}K_{fl}})\,\underline{\mathbf{e}}+\boldsymbol{h_{FL}}(t,\tilde{u})\\ f_{2}(t,\underline{\mathbf{e}},\tilde{u})\end{bmatrix}}_{\boldsymbol{f_{\chi}}(t,\boldsymbol{\chi})}+\underbrace{\begin{bmatrix}\boldsymbol{I_{3}}\\ \boldsymbol{0_{1\times 3}}\end{bmatrix}}_{\boldsymbol{B_{\chi}}}\boldsymbol{\xi_{F_{d}}}.$ (34)
+Lemma 1.
+The unperturbed closed-loop system $\boldsymbol{\dot{\chi}}=\boldsymbol{f_{\chi}}(t,\boldsymbol{\chi})$ ($\boldsymbol{\xi_{F_{d}}}=\boldsymbol{0}$) is Uniformly Globally Asymptotically Stable (UGAS).
+Proof.
+The system can be seen as the feedback interconnection of the following subsystems:
+[…]
+[…]
+as illustrated in Fig. 6.
+The following statements are true:
+  1. S1:
+    The unperturbed subsystem $\Sigma_{1}$ $(\tilde{u}=0)$ is Uniformly GES (UGES).
+  2. S2:
+    The origin of the subsystem $\Sigma_{2}$ is UGES.
+  3. S3:
+    There exist two scalar, positive-definite functions $V_{FL},W$ such that
+    $\frac{\partial V_{FL}(t,\underline{\mathbf{e}})}{\partial\underline{\mathbf{e}}}\boldsymbol{h_{FL}}(t,\tilde{u})=o(W(\underline{\mathbf{e}})),\;\forall(t,\underline{\mathbf{e}})\in[0,\infty)\times\mathbf{R}^{3}.$ (36)
+The validity of S1 directly follows from the fact that $\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}K_{fl}}$ is Hurwitz by design and hence, with $\tilde{u}=0$, the system $\dot{\underline{\mathbf{e}}}=(\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}K_{fl}})\,\underline{\mathbf{e}}$ has a UGES equilibrium at origin.
+The statement S2 is verified by the design of the estimator in (20), which under Conditions 1-3, guarantees exponential stability of the origin of the estimation error $\tilde{u}(t)$ at all times.
+Finally, consider the scalar, positive-definite Lyapunov function $V_{FL}=\underline{\mathbf{e}}^{T}P_{FL}\,\underline{\mathbf{e}}$, where $P_{FL}$ is the symmetric, positive-definite solution to the Lyapunov matrix equation
+$(\boldsymbol{P_{FL}A_{FL}}-\boldsymbol{B_{FL}K_{fl}})+(\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}K_{fl}})^{T}\boldsymbol{P_{FL}}=-\boldsymbol{I}$
+associated to the UGES origin of the unperturbed $\Sigma_{1}$. Then
+[…]
+$\leq\left\|2\boldsymbol{P_{FL}}\underline{\mathbf{e}}\right\|\,\left\|\boldsymbol{B_{FL}}\right\|\,L_{u}\,|\tilde{u}|\leq\underbrace{\left\|2\boldsymbol{P_{FL}}\right\|\,\left\|\boldsymbol{B_{FL}}\right\|\,L_{u}\,|\tilde{u}|_{\infty}}_{\eta>0}\,\left\|\underline{\mathbf{e}}\right\|$
+$=\eta\|\underline{\mathbf{e}}\|^{2}\triangleq W(\|\underline{\mathbf{e}}\|).$ (37)
+The last inequality implies $\frac{\partial V_{FL}(t,\underline{\mathbf{e}})}{\partial\underline{\mathbf{e}}}\boldsymbol{h_{FL}}(t,\tilde{u})=o(W(\underline{\mathbf{e}}))$, $\forall(t,\underline{\mathbf{e}})\in[0,\infty)\times\mathbf{R}^{3}$.
+Following the approach introduced in [loriaFeedbackCascadeinterconnectedSystems2008], the unperturbed closed-loop system (35) can be seen as a cascaded interconnection of $\Sigma_{1}$ and $\Sigma_{2}$, where the solutions $\underline{\mathbf{e}}(t;\underline{\mathbf{e}}_{\boldsymbol{0}})$ of $\Sigma_{1}$ for all initial conditions $\underline{\mathbf{e}}_{\boldsymbol{0}}$ can be thought of as time-varying parameters of $\Sigma_{2}$.
+Then since statements S1, S2, S3 are true, by Proposition 1 in [loriaFeedbackCascadeinterconnectedSystems2008] the solutions of the unperturbed closed-loop system are globally UGAS.
+∎
+Proposition 1.
+The solutions of the entire closed-loop system in Eq. (34) are Uniformly Ultimately Bounded (UUB).
+Proof.
+This follows directly from Lemma 1 and Lemma 9.3 in [khalilNonlinearSystems2002] since the bounded perturbation $\boldsymbol{\xi_{F_{d}}}$ is non-vanishing at the origin.
+Figure 6: Closed loop system $\boldsymbol{\dot{\chi}}=\boldsymbol{f_{\chi}}(t,\boldsymbol{\chi})+\boldsymbol{B_{\chi}}\boldsymbol{\xi_{F_{d}}}$.
 VI Case Study
@@
 The angular dynamics of the blade is described by:
-$J_{tot}\dot{v_{\beta}}=\left(A_{p}p_{p}-A_{r}p_{r}\right)r\sin\left(\psi\left(x_{c}\right)\right)-M_{fric}-M_{ext}$ (45)
+$J_{tot}\dot{v_{\beta}}=\overbrace{\left(A_{p}p_{p}-A_{r}p_{r}\right)r_{\beta}\sin\left(\psi\left(x_{c}\right)\right)}^{M_{h}}-M_{fric}-M_{ext}$ (38)
 $\dot{\beta}=v_{\beta},$
-where the equivalent mass $M$ also accounts for blade inertia.
-It is noticed that the equations match the dynamics of the cylinder dynamics presented in Eq.
+where the equivalent mass $M$ now also accounts for blade inertia.
+(38) matches the dynamics of the cylinder dynamics presented in (7).
+The presence of a cyclic component in the variable distance that relates mechanical moment to force is minor, as shown in Fig.
 7.
-The the presence of a cyclic component in the variable distance that relates mechanical moment to force is minor, as will be shown.
 Thus, the control design aims at tracking the angle while the cylinder dynamics represents a stable internal dynamics, under the assumption of having a stiff mechanical connection between cylinder and blade.
-Hence, there exists a static transformation relating $x_{c}$ to $\beta$ and with the pitch control error being the difference between pitch angle and a reference,
-$e=\beta-r,$ (46)
-the relative degree of the closed loop pitch control will be the same as for the closed loop cylinder piston position control loop analysed above.
-With relative degree three of both, the third derivative of the error is needed,
-$e^{(3)}=$ $-\frac{\dot{M}_{fric}+\dot{M}_{ext}}{J_{tot}}$ (47)
-$-\frac{lA_{p}^{2}B}{J_{tot}V_{p}}v_{c}\sin{\psi_{p}(x_{c})}-\frac{lA_{r}^{2}B}{J_{tot}V_{r}}v_{c}\sin{\psi_{p}(x_{c})}$
-$+\frac{lA_{p}B}{J_{tot}V_{p}}Q_{p}(u)\sin{\psi_{p}(x_{c})}+\frac{lA_{r}B}{J_{tot}V_{r}}Q_{r}(u)\sin{\psi_{p}(x_{c})}$
-$+\underbrace{\frac{l}{J_{tot}}F_{h}\cos(\psi_{p}(x_{c}))\frac{d\psi}{dx_{c}}v_{c}}_{\tau(x)}-r^{(3)}\,\triangleq\,\varpi.$ (48)
-With respect to the general case presented in Eq. 8, two variations are present:
-  • $1/M$ is replaced by $l\sin{\psi_{p}(x_{c})}/J_{tot}$.
+Hence, there exists a static transformation relating $x_{c}$ to $\beta$ and with the pitch tracking error $e_{\beta}$ being the difference between pitch angle and a reference,
+$e_{\beta}=\beta-r,$ (39)
+the relative degree of the closed loop pitch control will be the same as for the closed loop cylinder piston position control loop analyzed above.
+With relative degree three of both, the third derivative of the new tracking error is needed,
+$e_{\beta}^{(3)}=$ $-\frac{\dot{M}_{fric}+\dot{M}_{ext}}{J_{tot}}$ (40)
+$-\frac{r_{\beta}A_{p}^{2}B}{J_{tot}V_{p}}v_{c}\sin{\psi_{\beta}(x_{c})}-\frac{r_{\beta}A_{r}^{2}B}{J_{tot}V_{r}}v_{c}\sin{\psi_{\beta}(x_{c})}$
+$+\frac{r_{\beta}A_{p}B}{J_{tot}V_{p}}Q_{p}(u)\sin{\psi_{\beta}(x_{c})}+\frac{r_{\beta}A_{r}B}{J_{tot}V_{r}}Q_{r}(u)\sin{\psi_{\beta}(x_{c})}$
+$+\underbrace{\frac{r_{\beta}}{J_{tot}}F_{h}\cos(\psi_{\beta}(x_{c}))\frac{d\psi}{dx_{c}}v_{c}}_{\tau(x)}-r^{(3)}\,\triangleq\,v.$ (41)
+Figure 7: Momentum arm angle for verifying that constant assumption holds.
+With respect to the general case presented in (8), two variations are present:
+  • $1/M$ is replaced by $r_{\beta}\sin{\psi_{\beta}(x_{c})}/J_{tot}$.
   • An additional term $\tau(x)$, is introduced.
 These changes arise due to the variable-length moment arm in the updated dynamics.
-Figure 5: Momentum arm angle for verifying that constant assumption holds.
-However, Fig. 5 suggests that $\psi(x_{c})$ is approximately constant, and thus the following assumptions are made:
-  • $\psi(x_{c})\approx\textit{const}$
-  • $\frac{d\psi}{dx_{c}}\approx 0\rightarrow\tau(x)\approx 0$.
-VI-A Standard controllers for study case
-The control action for the standard cases are obtained by following the procedure described in Sec. III. The input gain is defined as
-$g=\frac{lA_{p}B}{J_{tot}V_{p}}K_{Qp}+\frac{lA_{r}B}{J_{tot}V_{r}}K_{Qr}.$ (49)
-The feedback linearization controller is derived as,
-[…]
-where the angle dependency is removed, since it only affects the $r^{(3)}$ term and the term associated with the gain, which is a degree of freedom anyway.
-The same applies for the sliding mode controller:
-$u_{ST}=\frac{1}{g}$ $\left[\frac{\dot{\hat{M}}_{d}}{J_{tot}}+\frac{lA_{p}^{2}B}{J_{tot}V_{p}}v_{c}+\frac{lA_{r}^{2}B}{J_{tot}V_{r}}v_{c}+r^{(3)}-\lambda_{2}\ddot{e}-\lambda_{1}\dot{e}\right.$ (51)
-$\,\,\left.-k_{1}\lfloor s\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s)d\tau\right],$
-with
-$\begin{bmatrix}\dot{e}\\ \ddot{e}\end{bmatrix}\approx\begin{bmatrix}v_{\beta}-\dot{r}\\ \displaystyle-\frac{\hat{M_{d}}}{J_{tot}}+\frac{l(A_{p}p_{p}-A_{r}p_{r})}{J_{tot}}-\ddot{r}\end{bmatrix}$ (52)
-and $\hat{M_{d}}$ the new disturbance being estimated.
-VI-B Mapping inversion based controllers for study case
-Following the same approach, the virtual control input in the proposed solutions are:
-$v_{FL}=\frac{\dot{\hat{M}}_{d}}{J_{tot}}+\frac{lA_{p}^{2}B}{J_{tot}V_{p}}v_{c}+\frac{lA_{r}^{2}B}{J_{tot}V_{r}}v_{c}+r^{(3)}-K_{FL}\underline{\mathbf{e}}$ (53)
-$v_{ST}=$ $\frac{\dot{\hat{M}}_{d}}{J_{tot}}+\frac{lA_{p}^{2}B}{J_{tot}V_{p}}v_{c}+\frac{lA_{r}^{2}B}{J_{tot}V_{r}}v_{c}+r^{(3)}-\lambda_{2}\ddot{e}-\lambda_{1}\dot{e}$ (54)
-$\,\,-k_{1}\lfloor s\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s(\tau))d\tau,$
-and the dynamical mapping inversion algorithm is,
+However, being $\psi_{\beta}(x_{c})$ is approximately constant, the following assumptions are made:
+  • $\psi_{\beta}(x_{c})\approx\textit{const}$
+  • $\frac{d\psi_{\beta}}{dx_{c}}\approx 0\rightarrow\tau(x)\approx 0$.
+VI-A Simulation Setup
+The designed control actions are tuned on a high-fidelity simulator, which includes the general dynamics presented in (1) and (38), and also encompass the control valve dynamics.
+The measured signals are subjected to noise, simulated to match the levels in the test rig.
+The velocities ($v_{c}$ and $v_{\beta}$) are obtained via numerical differentiation since they are not usually available as sensor signals in a real plant.
+The reference for the pitch angle and the wind load $M_{ext}$ are obtained from a wind turbine high-fidelity simulator based on OpenFAST, widely used in wind turbine research and development [Jonkman2013].
+VI-B Experimental setup
+The proposed algorithm is further tested on a full scale pitch system from a real wind turbine with a rated power of 3 MW.
+A picture of the setup is presented in Fig.
+8, and further details can be found in [pedersenOnlineLoadLeakage2025].
+Figure 8: Full-scale experimental setup from a wind turbine: $104\,\mathrm{kg}$ hydraulic cylinder with a rated force of $150\,\mathrm{kN}$, blade mount and pitch bearing with an inertia of $280\,\mathrm{kg\,m}^{2}$.
+The 80 m blade is not attached, but its inertia is emulated through the action of a load cylinder, which excites the setup in real time, and also includes the OpenFAST calculated blade moment.
+For practical reasons, the test rig does not include the wind turbine blade.
+However, an additional load cylinder is available to both replicate the wind load from OpenFAST and digitally reproduce the blade’s inertia.
+Fig.
+9 presents a detailed diagram of the test rig’s composition.
+In the real turbine, the total inertia is given by the inertia of the pitch test rig ($J_{tr}$) plus the inertia of the blade ($J_{b}$). One can write:
+$J_{tr}\dot{v_{\beta}}$ $=M_{h}-M_{fric}-M_{ext}-J_{b}\dot{v_{\beta}}$ (42)
+$\approx M_{h}-M_{fric}-\underbrace{\left(M_{ext}+J_{b}\,\ddot{r}\right)}_{M_{h,l}},$
+by assuming $r\approx\beta$, and avoiding double differentiating the noisy position measurement.
+Thus, the inertia of the blade is reproduced by injecting an additional mechanical torque by means of the load cylinder.
+The real-time implementation of the control action is performed using a Speedgoat real-time module for MATLAB/Simulink.
+VI-C Controllers Design
+The control action for the baseline feedback linearization controller and for the proposed algorithm are derived by adapting the methods described in Sec. III and Sec. IV, respectively.
+VI-C1 Standard controller
+The control action for the standard case is obtained by defining the input gain:
+$g=\frac{r_{\beta}A_{p}B}{J_{tot}V_{p}}K_{Qp}+\frac{r_{\beta}A_{r}B}{J_{tot}V_{r}}K_{Qr}.$ (43)
+Finally, the feedback linearization controller is derived as,
+$\begin{aligned} u&=\frac{1}{g}\left[\frac{r_{\beta}A_{p}^{2}B}{J_{tot}V_{p}}v_{c}+\frac{r_{\beta}A_{r}^{2}B}{J_{tot}V_{r}}v_{c}+\frac{\dot{\hat{M}}_{d}/J_{tot}+r^{(3)}-K_{FL}\underline{\mathbf{e}}_{\beta}}{\sin({\psi_{\beta}(x_{c})})}\right]\\ &\approx\frac{1}{g}\left[\frac{r_{\beta}A_{p}^{2}B}{J_{tot}V_{p}}v_{c}+\frac{r_{\beta}A_{r}^{2}B}{J_{tot}V_{r}}v_{c}+\frac{\dot{\hat{M}}_{d}}{J_{tot}}+r^{(3)}-K_{FL}\underline{\mathbf{e}}_{\beta}\right]\end{aligned},$ (44)
+where $\hat{M_{d}}$ is available from the load cylinder, and the $\psi_{\beta}$ angle dependency is removed since it only affects the $r^{(3)}$ term and the term associated with the gain, which is a degree of freedom anyway.
+The vector $\underline{\mathbf{e}}_{\beta}$ contains the tracking error $e_{\beta}$ and its derivatives, similarly to (10).
+Figure 9: Schematics of the experimental setup with pitch control cylinder right, generating $F_{h}$, and load cylinder left, generating $F_{h,L}$ in real time.
+VI-C2 Mapping inversion based controller
+Following a similar approach, the virtual control input in the proposed solution is:
+$v=\frac{\dot{\hat{M}}_{d}}{J_{tot}}+\frac{r_{\beta}A_{p}^{2}B}{J_{tot}V_{p}}v_{c}+\frac{r_{\beta}A_{r}^{2}B}{J_{tot}V_{r}}v_{c}+r^{(3)}-K_{FL}\underline{\mathbf{e}}_{\beta},$ (45)
+and the dynamical mapping inversion algorithm is
 $\dot{u}=\gamma\,\mu(t,\boldsymbol{x},u)\big(v-\rho(t,\boldsymbol{x},u)\big)$ (46)
-$\rho(t,\boldsymbol{x},u)=\frac{lA_{p}B}{J_{tot}V_{p}}Q_{p}(p_{p},p_{r},u)+\frac{lA_{r}B}{J_{tot}V_{r}}Q_{rv}(p_{r},u)$
+$\rho(t,\boldsymbol{x},u)=\frac{r_{\beta}A_{p}B}{J_{tot}V_{p}}Q_{p}(p_{p},p_{r},u)+\frac{r_{\beta}A_{r}B}{J_{tot}V_{r}}Q_{rv}(p_{r},u)$
 $\mu(t,\boldsymbol{x},u)=\frac{\partial K}{\partial u}\left(\frac{A_{p}B}{MV_{p}}\sqrt{\Delta_{pp}(u)}+\frac{A_{r}B}{MV_{r}}\sqrt{\Delta_{pr}(u)}\right),$
-where $K(u)$ is defined according to Eq. 31 as
-$K(u)=10^{-7}\left(-23.67u^{5}+40.73u^{3}+1.97u\right)\;.$ (56)
-VII Simulation Results
-The designed control actions are tested on a high-fidelity simulator, which includes the general dynamics presented in Eqs.
-1,45, and also encompass the control valve dynamics.
-The measured signals are subjected to noise, simulated at realistic levels.
-The velocities ($v_{c}$ and $v_{\beta}$) are obtained via numerical differentiation since they are not usually available as sensor signals in a real plant.
-The pitch angle reference and the disturbance $M_{ext}$ are obtained from a high-fidelity simulator for a wind turbine based on OpenFAST that is widely used in wind turbine research and development.
-Figure 6: Tracking performance of the Feedback linearization based controller, with and without applying the dynamical mapping inversion.
+with $K(u)$ defined according to (25) for the case of the real deteriorated valve, and $u$ according to (24).
+VII Results
+The controllers are tested with the same tuning both in simulation and in the experimental setup.
+Furthermore, the same gains are chosen for the baseline solution and the presented algorithm, with the latter having an additional tuning parameter $\gamma$ for regulating the adaptation rate in the dynamic inversion.
+Different solutions for different values for $\gamma$ are presented to analyze its impact.
 VII-A Tracking performance
-The tracking performance of the feedback linearization controller is presented in Fig.
-6, and in Fig.
-7 in a zoomed version.
-Three controllers are presented, for comparing the standard Feedback Linearizatrion (FL) solution with the one where the dynamic mapping inversion is applied (MI-FL) under the same tuning gains, and by overtuning (OT) the standard solution to be able to match the performance of the MI-FL controller.
-While both controllers are showing good tracking performance, the MI-FL (for a given tuning) shows a smaller error over the simulation window.
-By increasing the gains, it is possible to match its performance, with the consequence of getting an increase of unwanted chattering in the input.
+The experimental recordings for the two controllers are presented in Fig. 10, and in Fig. 11 in a zoomed version.
+Both the baseline FL and the proposed algorithm are able to track the reference with an error which is always smaller than 0.01 rad.
+Around 180 seconds tracking is temporary lost or both solutions, due to limitations in the common supply pressure shared by both hydraulic systems.
+Both controllers saturate and are then able to track the reference back as soon as the required force (hence flow) decreases for both.
+A quantitative comparison between the two solutions, discarding the interval of saturation, is presented later.
+Figure 10: Experimental tracking performance of the Feedback linearization based controller, with and without applying the dynamical mapping inversion.
 Figure 11: Zoomed on regions of interest for showing the tracking properties under different conditions.
-The zoomed plot gives a visualization of the input behaviour $u$, and of the valve position $x_{v}$.
-The MI-FL controller shows a reduced chatter in the regions where the pitch is varying in amplituide, while it shows a more chattery behaviour when the pitch is constant or slowly varying.
-Furthermore, the signal quality in $u$ and $x_{v}$ are equivalent for the MI solution, while part of the noise is filtered out by the valve dynamics in the standard solution, since the controller doesn’t have knowledge of the valve dynamics, nor of valve limitations.
-The results for the ST-SMC are stated in Fig. 8.
-Figure 8: Tracking performance of the Supertwisting sliding mode controllers, with and without applying dynamical mapping inversion.
-The tracking performances are more similar between the two solutions with respect to the FL controller.
-The OT version doesn’t improve the tracking performance although leading to a more chattery input.
-The disturbance rate that the controller needs to handle, is shown in Fig.
-9, where it is evident how higher gains lead to a larger amplitude in the command signal.
-Figure 9: Perturbation derivative of sliding mode dynamics under different STSMC.
-The unexpected dependency of the disturbance rate on the tuning gains is justified by the use of an observer for estimating the disturbance.
-Higher gains lead to a higher error in the estimation, which is affecting the overall control performance.
-The performance comparison of the observer with the different controllers is stated in Fig.
-10.
-The tuning doesn’t coincide with the bounds since higher gains are reflected with higher disturbance estimation errors and, thus, higher disturbance derivative bounds.
-The STSMC is not capable of ensuring convergence on the origin at any time instant, but boundedness is guaranteed, according to [24].
+However, from the zoomed plot it is already possible to appreciate some differences between the two solutions.
+The control input $u$ displays reduced chattering, especially in the region where no pitching action is present.
+The control action appears smoother also in the region of active pitching, even if it also contains more similar oscillations with respect to the baseline solution.
+Such difference is smoothened when analyzing the actual spool’s position $x_{v}$.
+In fact, the valve performs a natural filtering action according to its bandwidth.
+The signal quality in $u$ and $x_{v}$ are equivalent for the MI solution, while part of the noise is filtered out by the valve dynamics in the standard solution, since the controller doesn’t have knowledge of the valve dynamics, nor of valve limitations.
+Remark 4.
+The spikes in the valve position can be originated by the internal closed-loop or by the internal position sensor.
 VII-B Mapping inversion algorithm
 The tracking performance of the mapping inversion algorithm is presented in Fig.
-11.
-$v$ and $\rho$ are the signal defined in Eq.
-55, while $v_{sim}$ is the equivalent of the $v$ signal which is obtained in simulation, hence based on the real states of the system and not on the noisy measurements.
+12, both for the case of simulation and the experiment.
+The slight mismatch between the two can be related to model uncertainty, and to the fact that the digitally reproduced inertia is an equivalent load which wouldn’t be present in the real turbine.
+Its effect on the total load $M_{h,l}$ is the addition of a high-frequency component which causes low-amplitude high-frequency oscillations in the signals, reflected in a virtual input displaying larger peaks.
+$v$ and $\rho$ are the signal defined in (46), while $v_{sim}$ is the equivalent of the $v$ signal which is obtained in simulation, hence based on the real states of the system and not on the noisy measurements.
 The algorithm is not able to perfectly reconstruct $v$, because of the presence of noise in the signals.
+Figure 12: Dynamical mapping inversion algorithm: Simulation with correct virtual input reconstruction, and experiment.
 However, it is shown how it is possible to determine a proper tuning such that it is possible to reconstruct the actual virtual input in the plant ($v_{sim}$) and filtering out the noise present in the measurements.
-Figure 10: Wind load disturbance estimation and comparison with LPF hydraulic force.
-Thus, there is a natural filtering action which can be approximate with a simple first order low-pass filter in order to determine an indication of the bandwidth of such filtering action, in comparison to the bandwidth of the valve which represents a physical limitation in what the control action can deliver on the real system.
-The comparison is presented in Fig.
-12.
-Figure 11: Dynamical mapping inversion algorithm capability of tracking the required virtual controller, with a filtering action that allows to reconstruct the actual signal value in simulation
-The cutoff frequency of the MI algorithm is lower than the valve’s bandwidth, meaning the valve only needs to follow trajectories it can actually withstand.
+Thus, there is a natural filtering action which can be approximated with a simple first order low-pass filter in order to determine an indication of the bandwidth of such filtering action, in comparison to the bandwidth of the valve which represents a physical limitation in what the control action can deliver on the real system.
+The comparison, for $\gamma$ values of $(5,\,10,\,20)\times 10^{-4}$, is presented in Fig.
+13.
+The cutoff frequency of the MI algorithm is always lower than the valve’s bandwidth, meaning the valve only needs to follow trajectories it can actually withstand.
 With respect to the standard solution, where the filtering was occurring in the valve directly, in this new solution is the algorithm accounting for the valve limitations and filtering the input signal.
 This result support the similar signal quality in $u$ and $x_{v}$ in the MI solution that was found in Fig.
-7.
-Figure 12: Bode plots of valve dynamics and mapping inversion filtering effect.
+11.
 Remark 5.
 The tuning presented shows some room with respect to the maximum bandwidth that can be achieved by the valve.
-However, a faster tuning is reflected in a noisier reconstruction of $v_{sim}$ from the noisy measurements, hence more chattery $u$ and $x_{v}$.
+However, a faster tuning is reflected in a noisier reconstruction of $v_{sim}$ from the noisy measurements, leading to instability.
 The slew rate limitation of the valve may also play a role.
+Figure 13: Bode plots of valve dynamics and mapping inversion filtering effect for different adaptation gains.
 VII-C Controllers comparison
@@
 The same applies for determining $x_{v,f}$.
-An example of a $u_{f}$ signal is illustrated in Fig.
-13.
+The case of $u_{f}$, for the same moving average filter applied to both solutions, is illustrated in Fig.
+14.
 The following metrics are defined for the purpose:
-  • $\text{RMS}(u_{f})=\sqrt{\frac{1}{N}\sum_{i=1}^{N}u_{fi}^{2}}$
-  • $\text{RMS}(x_{v,f})=\sqrt{\frac{1}{N}\sum_{i=1}^{N}x_{v,fi}^{2}}$
-  • $\text{MAX}(e)=\max\limits_{0\leq t\leq T}e(t)$
-  • $\text{RMS}(e)=\sqrt{\frac{1}{N}\sum_{i=1}^{N}e_{i}^{2}}$
-  • $\text{QWAE}(e)=\sqrt{\frac{\sum_{i=1}^{N}e_{i}^{2}|e_{i}|}{\sum_{i=1}^{N}e_{i}^{2}}}$
+  • $\sqrt{\frac{\sum_{i=1}^{N}e_{\beta,i}^{2}|e_{\beta,i}|}{\sum_{i=1}^{N}e_{\beta,i}^{2}}}$$\text{RMS}(u_{f})=\sqrt{\frac{1}{N}\sum_{i=1}^{N}u_{fi}^{2}}$
+  • $\sqrt{\frac{\sum_{i=1}^{N}e_{\beta,i}^{2}|e_{\beta,i}|}{\sum_{i=1}^{N}e_{\beta,i}^{2}}}$$\text{RMS}(x_{v,f})=\sqrt{\frac{1}{N}\sum_{i=1}^{N}x_{v,fi}^{2}}$
+  • $\sqrt{\frac{\sum_{i=1}^{N}e_{\beta,i}^{2}|e_{\beta,i}|}{\sum_{i=1}^{N}e_{\beta,i}^{2}}}$$\text{MAX}(e_{\beta})=\max_{0\leq t\leq T}|e_{\beta}(t)|$
+  • $\sqrt{\frac{\sum_{i=1}^{N}e_{\beta,i}^{2}|e_{\beta,i}|}{\sum_{i=1}^{N}e_{\beta,i}^{2}}}$$\text{RMS}(e_{\beta})=\sqrt{\frac{1}{N}\sum_{i=1}^{N}e_{\beta,i}^{2}}$
+  • $\text{QWAE}(e_{\beta})=\sqrt{\frac{\sum_{i=1}^{N}e_{\beta,i}^{2}|e_{\beta,i}|}{\sum_{i=1}^{N}e_{\beta,i}^{2}}}$
+  • $\sqrt{\frac{\sum_{i=1}^{N}e_{\beta,i}^{2}|e_{\beta,i}|}{\sum_{i=1}^{N}e_{\beta,i}^{2}}}$$\text{ISE}(e_{\beta})=\int_{0}^{T}e_{\beta}^{2}(t)\,dt$
+where $T$ is the simulation time and the resulting values are normalized for having a clearer comparison.
+The QWAE (Quadratically Weighted Average Error) metric is used to penalize larger errors more heavily, while the ISE (Integral Squared Error) metric measures the persistence of errors.
+Finally, the time window where all the controllers saturate is removed from the evaluation, since it is not representative of the real turbine operation.
+Figure 14: Input signal $u$ and moving average $\bar{u}$ (upper), and the difference $u-\bar{u}$ (lower).
+Left column plots are with FL alone, right column is with the MI combined with FL.
+Chattering is significantly reduced with inclusion of MI.
+The experimental results for both controllers are presented in Fig.
+15, displaying how the solution proposed in the article outperforms the baseline for the same controller tuning and with the adaptation gain $\gamma=10\times 10^{-4}$, which corresponds to the tuning case $MIFL,10$ in Fig.
+16.
+The adaptive law was implemented using forward Euler integration, with a sampling time of 1ms.
+Figure 15: Experimental results comparing the baseline algorithm Fig.
+2 with the suggested MI approach Fig.
+3 .
+Blue bars of the baseline control are normalized to 1, and green bars (scaled accordingly) show the quality measures for the MIFL solution, relative to that of the baseline approach (FL).
+Introduction of the the mapping inversion has very significant benefits.
+Normalization value presented on top of each group of bars.
 […]
-where $T$ is the simulation time and the resulting values are normalized for having a clearer comparison.
-Figure 13: Moving average removal for deriving a metrics associated with signal chatter.
-The QWAE (Quadratically Weighted Average Error) metric is used to penalize larger errors more heavily, while the ISE (Integral Squared Error) metric measures the persistence of errors.
-The result for the system running with the different controllers is shown in Fig. 14.
-The solution proposed in the article performs better than standard one with the same tuning and with higher gains.
-The only exceptions are the cases of the $RMS(e)$ and the $ISE(e)$, where the OT ST-SMC displays a slightly smaller error, at the price of having much worse input performance.
-For a fixed tuning, the error with the MI is lower than when applying the respective standard algorithm.
-The input signal is significantly improved by the filtering action of the MI, while the improvement in the actual valve position becomes less significant thanks to the filtering action of the valve itself.
-Still, the solution that leverages the MI shows the cleanest input over a window where both fast varying and constant references are required.
-The amplitude of $\text{RMS}(u_{f})$ and $\text{RMS}(x_{v,f})$ are equivalent in the MI case.
-Figure 14: Bar plots for evaluating performance of the different controllers according to different metrics, associated with control input and tracking error.
+[…]
+Figure 16: Experimental results comparing different tunings for the input inversion algorithm.
+Higher values of $\gamma$ are reflected in smaller error indexes, but larger chattering amplitude.
+Normalization values on top of each bar group is added to make comparison possible with Fig.15.
+The results for three sets of tuning parameters of the inversion algorithm, see Fig.
+13, are presented in Fig.
+16.
+Except for the maximum error metrics, all the others share the expected trend: error values decrease with higher values of $\gamma$, while input chatter values increase.
+The normalization value presented on top of each bar group is always smaller than the respective one presented in Fig.
+15, meaning that the solution with the MI outperforms the baseline FL for all the tuning cases of the MI.
+The experimental results hence demonstrate that the MI approach, with the $\gamma$ tuning described, provides an easily implementable and effective solution that can obtain satisfactory tracking of an electro hydraulic actuator while obtaining significant reduction in wear caused by reduction in EHA micro movements and simultaneously by reduced energy in unwanted valve fluctuations.
 VIII Conclusion
-This article presented a solution to eliminate the formation of an algebraic loop when designing some conventional nonlinear controllers.
-This was achieved by design of a virtual input and dynamically deriving a control action, which emulated the desired behaviour, while control valve trajectory commands were avoided if they would exceed the specified admissible range for valve behaviours.
-The effectiveness of the proposed solution was tested in high-fidelity simulation for the industrial application of wind turbines’ blade pitch control, showing a general improvement in tracking performance and an essential reduction of input chatter, especially in the regions where the pitch reference is highly varying.
+This article presented a solution to eliminate the formation of an algebraic loop when designing conventional nonlinear controllers.
+This was achieved by design of a virtual input and dynamically deriving a control action, which emulated the desired behavior, while control valve trajectory commands were avoided if they would exceed the specified admissible range for valve operations.
+The effectiveness of the proposed solution was tested in high-fidelity simulation and experimentally for the industrial application of wind turbines’ blade pitch control, showing a general improvement in tracking performance and an essential reduction of input chatter, especially in the regions where the pitch reference is highly varying.
 The dynamic inversion solution enabled room for flexibility and allows for future progress with other nonlinear parameter estimation algorithms.
-A further step towards industrial implementation will be to derive the discrete-time version of the algorithm.
+A further step towards industrial implementation will be to derive the implicit discrete-time version of the algorithm.
 Acknowledgements
@@
 The authors much appreciate this support.
-[Feedback interconnected systems derivations] Let $\delta_{\underline{\mathbf{e}}}$ be the error associated with the calculation of $\ddot{e}$.
-The closed-loop system with the FL controller in the form of $\Sigma_{FL}$ is obtained as:
-$\dot{\underline{\mathbf{e}}}=$ $\boldsymbol{A_{FL}}\underline{\mathbf{e}}+\boldsymbol{B_{FL}}(v^{*}-\tilde{v})$ (57)
-$=$ $\boldsymbol{A_{FL}}\underline{\mathbf{e}}+\boldsymbol{B_{FL}}\left(-K_{FL}(\underline{\mathbf{e}}-\delta_{\underline{\mathbf{e}}})-\frac{\dot{\tilde{F}}_{d}}{M}-\tilde{v}\right)$
-$=$ $(\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}}K_{fl})\,\underline{\mathbf{e}}+\boldsymbol{B_{FL}}\left(K_{FL}\begin{bmatrix}0\\ -\frac{\tilde{F}_{d}}{M}\end{bmatrix}-\frac{\dot{\tilde{F}}_{d}}{M}\right)$
-$-\boldsymbol{B_{FL}}\,\left[\rho(\tilde{u})-\rho(u^{*}-\tilde{u})\right]$
-$=$ $(\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}}K_{fl})\,\underline{\mathbf{e}}+\boldsymbol{\xi_{F_{d}}}+\boldsymbol{h_{FL}}(t,\tilde{u}).$
-The unperturbed system is $\boldsymbol{h_{FL}}(t,0)=\boldsymbol{0}$ and the remaining terms describe the closed-loop dynamics with the standard design.
-Except for the bounded term $\boldsymbol{\xi_{F_{d}}}$, the system is exponentially stable with a convergence rate determined by the eigenvalues of $\boldsymbol{A_{FL}}-\boldsymbol{B_{FL}}K_{fl}$.
-The bounded term acts as a disturbance that causes the system to be UUGB with sufficiently high gains.
-Let
-$\tilde{s}=s-s^{*},$ (58)
-where $s$ is the actual value of the sliding variable and $s^{*}$ is calculated by differentiating the error with a robust exact differentiator, like the one in [25].
-The sliding variable dynamics in closed loop reads:
-$\dot{s}=$ $-k_{1}\lfloor s^{*}\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s^{*}(\tau))d\tau$ (59)
-$-k_{1}\lfloor\tilde{s}\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(\tilde{s}(\tau))d\tau-\frac{\dot{\tilde{F}}_{d}}{M}-\lambda_{2}\frac{\tilde{F}_{d}}{M}$
-$-\rho(u^{*}-\tilde{u})$
-$=$ $-k_{1}\lfloor s^{*}\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s^{*}(\tau))d\tau-\Delta(t)-\rho(u^{*}-\tilde{u}),$
-where all the terms in $\Delta$ are bounded: $\tilde{s}$ is bounded and with zero mean, since the reconstruction error with the robust exact differentiator is only related to noise, and the disturbance estimations are bounded (proof in Sec.
-III-C1).
-For $s\neq 0$ one can define [22]
-$\zeta_{1}\triangleq|s^{*}|^{1/2}sgn(s^{*})$ (60)
-[…]
-$\boldsymbol{\zeta}\triangleq\begin{bmatrix}\zeta_{1}&\zeta_{2}\end{bmatrix}^{T},$
-and rewrite the closed-loop system as:
-$\boldsymbol{\dot{\zeta}}=\frac{1}{|\zeta_{1}|}\begin{bmatrix}-\displaystyle\frac{1}{2}k_{1}&1\\ -k_{2}&0\end{bmatrix}\boldsymbol{\zeta}+\begin{bmatrix}0\\ -d\end{bmatrix}+\underbrace{\begin{bmatrix}-\displaystyle\frac{\rho(u-\tilde{u})}{|\zeta_{1}|}\\ 0\end{bmatrix}}_{\displaystyle\boldsymbol{h}(t,\tilde{u})},$ (61)
-with $k_{1},k_{2}$ selected to fulfil the conditions in [26].
-Since $h(t,0)=0$, the unperturbed system has a finite stable equilibrium point at the origin.
-This implies (see [22]) the existence of a diffeomorphism $T:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ such that for $\boldsymbol{\vartheta}=T(\boldsymbol{\zeta})$ and
-$\boldsymbol{\dot{\vartheta}}=\frac{\partial T}{\partial\boldsymbol{\zeta}}\left\{\frac{1}{|\zeta_{1}|}\begin{bmatrix}-\frac{k_{1}}{2}&1\\ -k_{2}&0\end{bmatrix}T^{-1}(\boldsymbol{\vartheta})-\begin{bmatrix}0\\ d\end{bmatrix}\right\}-\frac{\partial T}{\partial\boldsymbol{\zeta}}\begin{bmatrix}\displaystyle\frac{1}{|\zeta_{1}|}\\ 0\end{bmatrix}\rho(\tilde{u}-\hat{u}),$ (62)
-$\exists V_{ST}=\boldsymbol{\vartheta}^{T}P_{ST}\,\boldsymbol{\vartheta}$ such that $V_{2}\leq-c||\boldsymbol{\vartheta}||\triangleq-W_{ST}(||\boldsymbol{\vartheta}||)$.
+The authors would also like to thank Prof.
+H.
+C.
+Pedersen, as well as postdoctoral researchers M.
+Bhola and G.
+Wrat, and the students at Aalborg University, for their valuable assistance in conducting the experiments.